### v1

# A Machine Learning and Finite Element Framework for Inverse Elliptic PDEs via Dirichlet-to-Neumann Mapping

Abstract
Inverse problems for partial differential equations (PDEs) are crucial in numerous applications such as geophysics, biomedical imaging, and material science, where unknown physical properties must be inferred from indirect measurements. In this work, we address the inverse problem for elliptic PDEs by leveraging the Dirichlet-to-Neumann (DtN) map, which captures the relationship between boundary inputs and flux responses. Thus, this approach enables to solve the inverse problem that seeks the material properties inside the domain by utilizing the boundary data. Our framework employs an unsupervised machine learning algorithm that integrates a finite element method (FEM) in the inner loop for the forward problem, ensuring high accuracy. Moreover our approach is flexible to utilize partial observations of the boundary data, which is often the case in real-world scenarios. By incorporating carefully designed loss functions that accommodate discontinuities, the method refines coefficient reconstructions iteratively. This combined FEM and machine learning approach offers a robust, accurate solution strategy for a broad range of inverse problems, enabling improved estimation of critical parameters in applications from medical diagnostics to subsurface exploration.
keywords:
inverse problem, elliptic PDEs, EIT, variable coefficient, boundary condition mapping
[…]
1 Introduction
Inverse problems for partial differential equations (PDEs) are a fundamental area of study in applied and computational mathematics, with wide-ranging applications in fields such as geophysics [1], biomedical imaging [2, 3], and material science. These problems center on deducing unknown coefficients—often representing physical properties such as conductivity, diffusivity, or elasticity—from observed data [4, 5]. For instance, in medical imaging techniques such as Electrical Impedance Tomography (EIT), recovering these hidden material properties from indirect measurements enables visualization of conductivity distributions inside the human body, thereby enhancing diagnostic capabilities [6, 7, 8]. Similarly, in geophysical contexts, inferring subsurface properties such as permeability from surface measurements is critical for resource exploration and environmental monitoring [9, 10].
To address these challenges, we focus on the inverse problem for an elliptic equation with variable, potentially discontinuous coefficients. The main novelty of our approach is the utilization of the Dirichlet-to-Neumann (DtN) map—a mathematical construct that encapsulates the relationship between prescribed boundary inputs and their resulting flux responses [11]. This mapping framework is especially advantageous because it allows for the reconstruction of the interior characteristics of a domain using solely boundary measurements. In scenarios such as EIT or subsurface exploration, where sensors are predominantly deployed on the boundary, this capability eliminates the need for invasive internal data collection while still providing detailed insight into the material properties within the domain. See Figure 1 for the illustration.
[…]
Figure 1: Illustration of the problem. The dashed line indicates the partially observed boundary measurement.
In addition, our approach is flexible to utilize partial observations of the boundary data, which is often the case in real-world scenarios. The presented numerical examples further illustrate the capabilities and robustness of the proposed algorithm, demonstrating its effectiveness across a range of challenging inverse problems.
The proposed computational framework further distinguishes itself by integrating an unsupervised machine learning algorithm with finite element methods (FEM) used in the inner loop for solving the forward problem. This combination not only achieves high accuracy in modeling the physical processes but also enables iterative refinement of the coefficient reconstructions. The carefully designed loss functions, tailored to handle discontinuities effectively, enhance the overall fidelity of the inverse solution, making our approach robust and versatile. Ultimately, this framework offers a precise and reliable estimation of critical parameters, paving the way for significant advancements in applications ranging from non-invasive medical diagnostics to environmental and subsurface monitoring.
The rest of the paper is organized as follows. First, the governing system and the DtN mapping are introduced in Section 2. The overall computational framework including the loss functions is discussed in Section 3. Finally, in Section 4, several numerical experiments to illustrate the capabilities and the performance of our algorithm are presented.
2 Governing System
In this paper, we consider the following equation
$\left\{\begin{array}[]{rll}-\nabla\cdot(k({\mathbf{x}})\nabla p({\mathbf{x}}))&=f({\mathbf{x}})&{\mathbf{x}}\in\Omega,\\ p({\mathbf{x}})&=g({\mathbf{x}})&{\mathbf{x}}\in\partial\Omega,\end{array}\right.$ (1)
where $\Omega$ is a bounded domain in $\mathbb{R}^{2}$, $p:\overline{\Omega}\rightarrow\mathbb{R}$ is the solution function, $f$ is the source function, $g\in H^{1/2}(\partial\Omega)$ is the given Dirichlet boundary condition, and $k$ is a strictly positive coefficient defined in $\Omega$.
We consider an inverse problem associated with (1), where the goal is to reconstruct the coefficient $k$ from the boundary measurements. To achieve this, we employ deep learning techniques. A straightforward approach would be to use supervised learning, which requires access to exact coefficient values of $k({\mathbf{x}})$ for all ${\mathbf{x}}\in\Omega$. However, in most cases, obtaining complete knowledge of $k$ in $\Omega$ is infeasible. Thus, we propose an unsupervised learning framework that does not rely on exact coefficient values of $k({\mathbf{x}})$ but instead leverages available data to infer the coefficient.
The methodology and key concepts necessary for solving this problem are introduced in the following subsection.
2.1 Dirichlet-to-Neumann Map
In this subsection, we introduce the (Dirichlet-to-Neumann) DtN map, which is central to the reconstruction of $k$. This map is defined as follows:
$\Lambda_{k}:H^{1/2}(\partial\Omega)\to H^{-1/2}(\partial\Omega),$
which maps the Dirichlet data $g$ to the corresponding Neumann data $h_{g}=k\partial_{\nu}p_{g}|_{\partial\Omega}$. Here, $p_{g}$ is the solution of (1) for the given source function $f$ and Dirichlet condition $g$, and $\nu$ denotes the unit outward normal vector on $\partial\Omega$. The space $H^{r}(\partial\Omega)$, $r\in\mathbb{R}$, denotes the Sobolev space.
The reconstruction of the coefficient $k$ using the DtN map has been extensively studied in the context of Electrical Impedance Tomography (EIT) and the Calderón problem [11]. In particular, the case when $f=0$ has been comprehensively surveyed in [12]. Furthermore, additional studies have explored the uniqueness of $k$ under partially observed boundary data. Let $\Gamma\subset\partial\Omega$ be a non-empty open subset of the boundary. The following theorem proves the uniqueness of $k$.
Theorem A (Corollary 1.19 in [12])
Let $\Omega\subset\mathbb{R}^{2}$ be a bounded domain with a smooth boundary, and let $k_{1},k_{2}\in C^{4+\alpha}(\overline{\Omega})$ be non-vanishing functions for some $\alpha>0$. Assume that
$\Lambda_{k_{1}}(g)=\Lambda_{k_{2}}(g)\text{ on }\Gamma\text{ for all }g\in H^{1/2}(\partial\Omega),\text{ supp }g\subset\Gamma.$
Then, $k_{1}=k_{2}$ in $\Omega$.
Furthermore, if $\Gamma=\partial\Omega$, the uniqueness of $k$ can be established using data from the entire boundary.
Although the case $f\neq 0$ may seem easier because $f$ contains internal information, it remains an open problem. While the uniqueness of $k$ in the entire domain $\Omega$ is not guaranteed, its uniqueness on the boundary $\partial\Omega$ has been established [13].
Theorem B (Theorem 5.1.1 in [13])
Assume that $k_{1},k_{2}\in C^{m}(V)$, where $V$ is a neighborhood of a boundary point of $\Omega$.
[…]
If $k_{1},k_{2}\in C(\Omega)$ are two coefficients that are analytic in $\Omega$ and $\Lambda_{k_{1}}=\Lambda_{k_{2}}$, then $k_{1}=k_{2}$ on $\partial\Omega$ and by Strong Maximum Principle [14, Chapter 2], we have $k_{1}=k_{2}$ in $\Omega$.
Let $\omega$ be either $\Gamma$ or $\partial\Omega$. We refer to the case $\omega=\Gamma$ as the partially observed case, and $\omega=\partial\Omega$ as the fully observed setting. Instead of directly using the DtN map, which is usually difficult to obtain in real-world scenarios, we utilize Cauchy data—pairs of Dirichlet and Neumann boundary measurements.
$\left\{(g,h_{g}):g\in H^{1/2}(\partial\Omega),\text{ supp }g\subset\omega,h_{g}=k\partial_{\nu}p_{g}|_{\omega}\right\}.$
These pairs serve as practical substitutes for the DtN map in our setup. Again, in practice, it is not feasible to train on all functions $g\in H^{1/2}(\partial\Omega)$ with supp $g\subset\omega$. We consider the following problem.
Problem
Given a sufficiently large finite collection of Cauchy data:
$\left\{(g_{i},h_{g_{i}})_{i=1,\dots,N}:g_{i}\in H^{1/2}(\partial\Omega),\text{supp }g_{i}\subset\omega,h_{g_{i}}=k\partial_{\nu}p_{g_{i}}|_{\omega}\right\},$ (2)
the coefficient $k$ can be numerically reconstructed in $\Omega$.
Rather than directly using the values of $k$, we adopt an unsupervised learning framework where the model learns to reconstruct $k$ from Cauchy data data alone [15, 16]. Using this approach, we reconstruct $k$ based only on Cauchy data, which is particularly useful in practical setups where $k$ is not directly measurable.
3 Computational Algorithm
In this section, we present the numerical approach for solving Problem. First, in Subsection 3.1, we provide an overview of the proposed framework. We introduce the model for reconstructing the coefficient $k$ in Subsection 3.2. Next, in Subsection 3.3, we describe how to solve the forward problem using the model described in 3.2. Finally, in Subsection 3.4, we define the loss function used to train the model and describe the method for calculating its gradient to update the model parameters.
3.1 Overall Framework
In this subsection, we introduce the proposed computational framework for reconstructing the coefficient $k$.
Input $\bullet$ Domain points
$\overline{\Omega}$
$\bullet$ Neumann data
$h_{g}=k\partial_{\nu}p_{g}|_{\omega}$
Neural Network $\mathcal{N}$
$\bullet$ Multilayer Perceptron (MLP)
: Computation results: Data flow Loss Function
$\mathcal{L}\bigl{(}h_{g},\tilde{h}_{g}\bigr{)}=\bigl{\|}h_{g}(\cdot)-\tilde{h}_{g}(\cdot)\bigr{\|}^{2}_{L^{2}(\omega)}$
Output $\bullet$ Reconstruct Coefficients
$\{\tilde{k}({\mathbf{x}})\,|\,{\mathbf{x}}\!\!\in\!\!\overline{\Omega}\}$
Predicted Neumann
$\tilde{h}_{g}=\tilde{k}\,\partial_{\nu}\tilde{p}_{g}|_{\omega}$
Solving for the PDE (FEniCS) Input $\bullet$ Source function $f$
$\bullet$ Dirichlet condition $g$
$\tilde{p}_{g}({\mathbf{x}})=\sum_{j=1}^{M}p_{j}\,\phi_{j}({\mathbf{x}})$ Parameter UpdatesDeep LearningFinite Element Method
Figure 2: Framework for reconstructing $k$ using a neural network and FEM.
Figure 2 illustrates the process of reconstructing the coefficient $k$ using a neural network and FEM. The framework proceeds as follows: First, the domain points $\overline{\Omega}$ are input to the neural network $\mathcal{N}$, which outputs the reconstruction $\tilde{k}$ of $k$. The neural network $\mathcal{N}$ is designed as a multilayer perceptron (MLP) that reconstructs $k$ from the input domain points. Next, the reconstructed coefficient $\tilde{k}$, along with the Dirichlet condition $g$ and source function $f$, is used to solve the (1) via FEM with the FEniCS package. The resulting predicted solution is denoted as $\tilde{p}_{g}$, which is represented as a linear combination of basis functions $\{\phi_{i}\}_{i=1}^{M}$ in the finite-dimensional subspace $V_{h}$. These basis functions and the finite-dimensional subspace are further detailed in Subsection 3.3. Using $\tilde{k}$ and $\tilde{p}_{g}$, the predicted Neumann data $\tilde{k}\partial_{\nu}\tilde{p}_{g}|_{\omega}$ is computed. This predicted Neumann data is then compared with the given data $h_{g}$ to calculate the loss function. Finally, the model parameters of the neural network $\mathcal{N}$ are updated to minimize the loss function. By iterating through this process, the neural network $\mathcal{N}$ learns to reconstruct $k$ more accurately.
3.2 Model for Reconstructing the Coefficient $k$
In this subsection, we propose a model for reconstructing the coefficient $k$. The model is based on a MLP architecture, which leverages the Universal Approximation Theorem [17] as its theoretical foundation. This theorem ensures that neural networks, including MLPs, can approximate any continuous function to a desired level of accuracy.
The neural network $\mathcal{N}$ takes the spatial variable ${\mathbf{x}}$ as input and reconstructs the coefficient $k$ as $\tilde{k}$. Specifically, the reconstruction is expressed as $\tilde{k}({\mathbf{x}})=\mathcal{N}({\mathbf{x}};\boldsymbol{\theta})$, where $\boldsymbol{\theta}$ represents the learnable parameters of the model, such as weights and biases. The network consists of three hidden layers, each containing 50 hidden units. To enhance the ability of the model to capture periodic patterns and their derivatives, we employ the sine function as the activation function in the hidden layers. This choice enables the model to efficiently represent complex natural signals with high accuracy.
3.3 Forward Problem
In this subsection, we briefly describe the forward problem for the solution $p$ of (1) by utilizing FEM. Here, we employ the linear continuous Galerkin method, and it is implemented by FEniCS [18]. The problem assumes a given Dirichlet condition $g$ and coefficient $k$, and the problem (1) is reformulated into a weak form by integration and then approximating the solution in a finite-dimensional space.
In particular, let $V_{h}$ be a finite-dimensional subspace of $H^{1}(\Omega)$, and assume $p_{h},v_{h}\in V_{h}$, with $p_{h}$ representing the approximate solution and $v_{h}$ being the test function used in the weak formulation. Using the basis functions $\{\phi_{i}\}_{i=1}^{M}$ of $V_{h}$, the weak form can be approximated as:
$\int_{\Omega}k\nabla p_{h}\cdot\nabla v_{h}\,{\rm d}{\mathbf{x}}=\int_{\Omega}fv_{h}\,{\rm d}{\mathbf{x}}.$
Assuming $p_{h}=\sum_{i=1}^{M}P_{i}\phi_{i}$ and substituting $v_{h}=\phi_{j}$, the resulting system for each $j$-th term is:
$\sum_{i=1}^{M}P_{i}\int_{\Omega}k\nabla\phi_{i}\cdot\nabla\phi_{j}\,{\rm d}{\mathbf{x}}=\int_{\Omega}f\phi_{j}\,{\rm d}{\mathbf{x}}.$
This can be written in matrix form as:
$\mathbf{A}\mathbf{P}=\mathbf{b},$
where
$\mathbf{A}_{ij}=\int_{\Omega}k\nabla\phi_{i}\cdot\nabla\phi_{j}\,{\rm d}{\mathbf{x}},\quad\mathbf{P}=\begin{bmatrix}P_{1}\\ \vdots\\ P_{M}\end{bmatrix},\text{ and }\mathbf{b}=\begin{bmatrix}\int_{\Omega}f\phi_{1}\,{\rm d}{\mathbf{x}}\\ \vdots\\ \int_{\Omega}f\phi_{M}\,{\rm d}{\mathbf{x}}\end{bmatrix}.$
Here, $\mathbf{A}$ is known as the stiffness matrix, and $\mathbf{b}$ is the load vector. Finally, solving $\mathbf{A}\mathbf{P}=\mathbf{b}$ yields $\mathbf{P}$, which determines the coefficients for the basis functions, allowing us to obtain the approximate solution $p_{h}$.
3.4 Loss Functions
This subsection defines the loss function used to train the model introduced in Subsection 3.2, and describes how to compute its gradient with respect to the model parameters.
The loss function based on the Problem is defined as
$\mathcal{L}(\mathcal{D},\tilde{\mathcal{D}})=\sum_{i=1}^{N}\bigl{\|}h_{g_{i}}-\tilde{k}\partial_{\nu}\tilde{p}_{g_{i}}\bigr{\|}_{L^{2}(\omega)}^{2},$ (3)
where $\mathcal{D}=\{(g_{i},h_{g_{i}})_{i=1,\dots,N}\mid g_{i}\in H^{1/2}(\partial\Omega),\text{ supp }g_{i}\subset\omega,h_{g_{i}}=k\partial_{\nu}p_{g_{i}}|_{\omega}\}$, $\tilde{\mathcal{D}}=\{(g_{i},\tilde{k}\partial_{\nu}\tilde{p}_{g_{i}}|_{\omega})_{i=1,\dots,N}\mid g_{i}\in H^{1/2}(\partial\Omega),\text{ supp }g_{i}\subset\omega\}$, and $\|\cdot\|_{L^{2}(\omega)}$ denotes the $L^{2}$ norm on $\omega$. The loss function $\mathcal{L}$ is minimized by adjusting the model parameters $\boldsymbol{\theta}$. When the loss function reaches its minimum, we have $\mathcal{D}=\tilde{\mathcal{D}}$.
To optimize the model, we need the gradient of the loss function $\mathcal{L}$ with respect to the parameters $\boldsymbol{\theta}$. However, since this gradient cannot be computed directly, we apply the chain rule:
$\frac{\partial\mathcal{L}}{\partial\boldsymbol{\theta}}=\frac{\partial\mathcal{L}}{\partial k_{\text{fem}}}\cdot\frac{\partial\tilde{k}}{\partial\boldsymbol{\theta}}.$ (4)
[…]
4 Numerical Experiments
In this section, we present numerical experiments to validate and demonstrate the effectiveness of the proposed framework introduced in Section 3. The following setup is common across all examples. The entire dataset consists of 4,096 samples: 2,048 for training, 1,024 for validation, and 1,024 for testing. The neural network is trained for 100 epochs with a batch size of 1. The learning rate is set to 0.001 for the first 50 epochs and reduced to 0.0001 for the remaining 50 epochs to facilitate stable convergence and prevent overshooting during the later stages of training. Training is performed using PyTorch [19] and the ADAM optimizer [20], a variant of stochastic gradient descent. All experiments are conducted on an NVIDIA GeForce RTX 3090 GPU (24GB VRAM), an Intel Core i9-12900K CPU, and 62GB RAM. The system runs with CUDA 11.8 and NVIDIA driver version 522.06.
In the next subsections, we evaluate the performance of the proposed method in several cases:
4.1 Example 1: Constant Coefficient
In this example, we consider a constant coefficient $k$ to provide a baseline for evaluating model performance under simple conditions. Specifically, we set $k=1$. The computational domain is a unit square $\Omega=(0,1)\times(0,1)$ represented by a mesh with 66,049 points. The Dirichlet boundary condition is given by $g=\cos(a\pi x)\cos(b\pi y)|_{\partial\Omega}$, where, $a$ and $b$ are independently sampled from a uniform distribution over $[0,2]$. The source function is defined as $f(x,y)=(a^{2}+b^{2})\pi^{2}\cos(a\pi x)\cos(b\pi y)$.
To evaluate how effectively our method reconstructs the coefficient $\tilde{k}$, we employ two measures: the Neumann data loss and the relative error of $k$. The Neumann data loss, defined in (3) with $\omega=\partial\Omega$ (fully observed setting), measures the difference between the predicted and observed Neumann data. Meanwhile, the relative error provides a normalized measure of the difference between $\tilde{k}$ and the exact coefficient $k$:
$\|\tilde{k}\|_{\text{rel}}=\frac{\|k-\tilde{k}\|_{L^{2}(\overline{\Omega})}}{\|k\|_{L^{2}(\overline{\Omega})}},$ (5)
where $\|\cdot\|_{L^{2}(\overline{\Omega})}$ denotes the $L^{2}$ norm in $\overline{\Omega}$.
Figure 3 presents the Neumann data loss (left) and the relative error of $k$ (right) side by side on a logarithmic scale, where each point represents the averaged value over a 10 epoch interval. This averaging smooths out short-term fluctuations, providing a clearer view of the overall training dynamics.
Figure 3: (Left) Neumann training and validation loss (3) with $\omega=\partial\Omega$, and (Right) relative error of $k$ over epochs (log scale). Each point represents the averaged value over a 10 epoch interval.
As shown in the left plot of Figure 3, the training loss (blue dashed line) decreases sharply at the beginning of the training process and continues to decrease, reaching around $10^{-4}$ during 20–50 epochs. After 50 epoch, when the learning rate is reduced, the training loss continues to decrease and finally stabilizes around $10^{-6}$. The validation loss (red line) shows larger oscillations due to the smaller validation set size and the fact that it is computed once per training epoch, which increases its sensitivity to parameter updates. Despite this variability, the validation loss also experiences a significant decrease after 50 epoch, reaching approximately $10^{-8}$. These observations suggest that the model makes effective use of the Neumann boundary data, with the learning rate schedule playing a crucial role in refining the parameter updates. On the right, the relative error of $k$ follows a similar trend as the Neumann training loss (blue dashed line on the left). After a rapid initial decrease, it remains around $10^{-3}$ during 20-50 epochs, then undergoes a sharp decrease and stabilizes approximately $10^{-4}$. This simultaneous decrease in both metrics indicates that reducing the Neumann data loss effectively guides $\tilde{k}$ to the exact coefficient $k$.
These results support the claim in Problem that $k$ can be uniquely determined from Cauchy data.
(a) exact $k$
(b) reconstructed $\tilde{k}$
(c) absolute error $|k-\tilde{k}|$
Figure 4: Visualizations of the exact coefficient $k$, the reconstructed coefficient $\tilde{k}$, and the absolute error $|k-\tilde{k}|$.
Figure 4 visualizes the exact coefficient $k$, the reconstructed coefficient $\tilde{k}$, and the absolute error $|k-\tilde{k}|$, providing a clear assessment of the reconstruction accuracy. The results show that the proposed method yields a highly accurate approximation of $k$ with only minor deviations. This confirms the effectiveness of the approach in reconstructing constant coefficients with minimal error.
Up to this point, our analysis of the Neumann data loss and reconstruction error has demonstrated that the model can accurately reconstruct a constant coefficient. These numerical results confirm that training based on Neumann data plays a crucial role in optimizing the model. To provide a more intuitive understanding of how the boundary conditions operate in practice, we now examine the operation of the DtN map as illustrated in Figure 5.
[…]
Figure 5: Illustrations of the Dirichlet-to-Neumann (DtN) map $\Lambda$ for a 2D problem. The solution $p(x,y)$ is shown as a heat map. The Dirichlet boundary values $p(x,1)$ and $p(1,y)$ are mapped to their corresponding Neumann data via the DtN map.
The heat map on the left represents the solution $p(x,y)=\cos(0.3\pi x)\cos(0.6\pi y)$ over a rectangular domain. From this solution, Dirichlet data are obtained along the top and right boundaries, given by $g_{\text{Top}}(x)=p(x,1)$ and $g_{\text{Right}}(y)=p(1,y)$. The DtN map converts these Dirichlet data into the corresponding Neumann data, $h_{g_{\text{Top}}}(x)$ and $h_{g_{\text{Right}}}(y)$, as shown in the adjacent plots.
4.2 Example 1-2: Constant Coefficient with Partially Observed Data
In this example, we address the inverse problem of the partially observed data. In many real-world scenarios, measurements may be unavailable in certain regions due to sensor limitations or other constraints.
We consider the case of a constant coefficient $k=1$. The computational domain defined as the unit square $\Omega=(0,1)\times(0,1)$, represented by a mesh with 66,049 points. To simulate partial observations, we assume that measurements are unavailable in certain parts of the boundary, covering approximately 20% of $\partial\Omega$. Specifically, we define the region $\Gamma_{0}$, where measurements are unavailable, as follows: measurements are unavailable for $y\in[\frac{3}{5},\frac{4}{5}]$ on the left boundary ($x=0$), for $y\in[\frac{1}{5},\frac{2}{5}]$ on the right boundary ($x=1$), for $x\in[\frac{2}{5},\frac{3}{5}]$ on the bottom boundary ($y=0$), and for $x\in[\frac{3}{5},\frac{4}{5}]$ on the top boundary ($y=1$). In these regions, both the Dirichlet boundary condition and the source function are set to zero; that is, $g=0$ and $f=0$ on $\Gamma_{0}$. The remaining boundary, denoted by $\Gamma=\partial\Omega\setminus\Gamma_{0}$, is where measurements are available. On $\Gamma$, the Dirichlet boundary condition is defined as $g(x,y)=\cos(a\pi x)\cos(b\pi y)$, where $a$ and $b$ are independently sampled from a uniform distribution over $[0,2]$. Similarly, the source function is defined on $\Gamma$ as $f(x,y)=(a^{2}+b^{2})\pi^{2}\cos(a\pi x)\cos(b\pi y)$. The Neumann data is measured only on $\Gamma$, and the Neumann data loss function (3) is defined with $\omega=\Gamma$.
Figure 6: (Left) Boundary regions and (Right) Dirichlet boundary condition $g$.
Figure 6 illustrates the boundary region (left) and the corresponding Dirichlet boundary condition $g$ (right).
Figure 7: Relative error of $k$ over epochs (log scale) for both fully and partially observed settings. Each point represents the average over a 10-epoch interval.
(a) exact $k$
(b) reconstructed $\tilde{k}$
(c) absolute error $|k-\tilde{k}|$
Figure 8: Visualizations of the exact coefficient $k$, the reconstructed coefficient $\tilde{k}$, and the absolute error $|k-\tilde{k}|$ in the partially observed setting.
Figure 7 shows the relative error of $k$ on a logarithmic scale, comparing the cases of fully observed ($\omega=\partial\Omega$) and partially observed ($\omega=\Gamma$). Despite the limited observations in the partially observed setting, the relative error shows a convergence trend similar to that of the fully observed setting. This suggests that the proposed method remains robust even when only partial observations are available, achieving comparable accuracy. Figure 8 presents the exact coefficient $k$, the reconstructed coefficient $\tilde{k}$, and the absolute error $|k-\tilde{k}|$, providing a visual assessment of the reconstruction performance in the partially observed setting. The results indicate that the proposed method accurately reconstructs the coefficient even with partial boundary measurements, demonstrating its robustness in handling incomplete data.
4.3 Example 2: Spatially Varying Coefficient
In this example, we test the algorithm with a spatially varying function $k$. In this case, we set
$k(x,y)=0.9\sin\left(\frac{\pi}{3}(x+y+0.1)\right).$
All the other setups including the computational domain, number of mesh points, and Dirichlet boundary condition, are the same as the previous Example 1. The source function is defined as $f(x,y)=\nabla\cdot(k(x,y)\nabla(\cos(a\pi x)\cos(b\pi y)))$ by computing the divergence of the flux associated with the given function.
(a) exact $k$
(b) reconstructed $\tilde{k}$
(c) absolute error $|k-\tilde{k}|$
Figure 9: Illustrations of the exact coefficient $k$, the reconstructed coefficient $\tilde{k}$ in three dimensions, and the absolute error $|k-\tilde{k}|$ in a heat map.
Figure 9 illustrates the exact coefficient $k$, the reconstructed coefficient $\tilde{k}$ in a three-dimensional space, and their absolute error in a heat map, providing a comparative view of their structures. The reconstructed coefficient $\tilde{k}$ closely matches the exact $k$, effectively capturing its spatial variations, with the absolute error remaining at a low level.
To further analyze the reconstruction accuracy, Figure 10 presents slices of the coefficient $\tilde{k}$ and $k$ at the domain boundaries ($y=0$, $x=0$, $y=1$, and $x=1$) as well as interior slices along $y=x$ and $y=0.5$.
(a) y = 0
(b) x = 0
(c) y = 1
(d) x = 1
(e) y = x
(f) y = 0.5
Figure 10: Slices of the coefficient $\tilde{k}$ for the case of $k=0.9\sin\left(\dfrac{\pi}{3}(x+y+0.1)\right)$.
Figure 10 shows the stability of the reconstruction process by comparing both the boundary and interior regions. Over the entire domain, the reconstructed coefficient $\tilde{k}$ closely follows the behavior of $k$, showing strong agreement both at the boundaries and the interior. This consistency highlights the robustness of the method in accurately capturing spatial variations.
4.4 Example 2-2: Spatially Varying Coefficient with Partially Observed Data
In this example we consider the case of spatially varying coefficients on the partially observed data.
The coefficient $k$ is defined as $k(x,y)=0.9\sin(\frac{\pi}{3}(x+y+0.1))$. All other setups including the computational domain, number of mesh point, Dirichlet condition, and loss function are the same as the previous Example 1-2. The source function is defined as $f(x,y)=\nabla\cdot(k(x,y)\nabla(\cos(a\pi x)\cos(b\pi y)))$ by computing the divergence of the flux associated with the given function. This function is imposed on $\Gamma$, while $f$ is set to zero on $\Gamma_{0}$.
(a) exact $k$
(b) reconstructed $\tilde{k}$
(c) absolute error $|k-\tilde{k}|$
Figure 11: Visualizations of the exact coefficient $k$, the reconstructed coefficient $\tilde{k}$ in three dimensions, and the absolute error $|k-\tilde{k}|$ in a heat map in the partially observed setting.
Figure 11 presents the exact coefficient $k$, the reconstructed coefficient $\tilde{k}$ in three-dimensional space, and the absolute error $|k-\tilde{k}|$ in a heat map, providing a visual comparison of the reconstruction performance under partially observed setting. The results confirm the robustness of the proposed method in accurately reconstructing the coefficient even with limited boundary measurements.
Case Observed MSE MAE PSNR SSIM
Constant Fully $1.45\times 10^{-12}$ $1.05\times 10^{-6}$ 118.38 0.8346
Partially $1.38\times 10^{-8}$ $1.00\times 10^{-3}$ 78.61 0.9438
Spatially varying Fully $2.65\times 10^{-6}$ $1.50\times 10^{-3}$ 54.85 0.999996
Partially $2.47\times 10^{-7}$ $3.80\times 10^{-4}$ 65.16 0.999997
Table 1: Comparison of MSE, MAE, PSNR, and SSIM for the constant and spatially varying cases under fully and partially observed settings.
Table 1 presents the mean squared errors (MSE) for the results of Examples 1, 1-2, 2, and 2-2, comparing the fully and partially observed settings for both constant and spatially varying cases. In the constant case, the model achieves extremely low reconstruction error, with an MSE of $1.45\times 10^{-12}$ in the fully observed setting. Even in the partially observed setting, the MSE remains low at $1.37\times 10^{-8}$, indicating strong reconstruction performance despite incomplete boundary data. In the spatially varying case, the reconstruction is more challenging due to the spatial variability of the coefficient $k$. Nonetheless, the model achieves a low MSE of $2.65\times 10^{-6}$ in the fully observed setting, demonstrating its capability to approximate the target function accurately. Interestingly, the MSE in the partially observed setting is lower than in the fully observed setting. This may be attributed to the fact that both models converge to highly accurate solutions with very small errors, where minor numerical variations can lead the partially observed setting to yield a slightly lower MSE.
4.5 Example 3: Phantom Coefficient with a Zero Source Function
In this subsection, we consider the phantom coefficient, which is used in many medical imaging applications such as EIT [6, 7, 8] and PAT.
In this case, we only consider the fully observed setting, where $\omega=\partial\Omega$. The source function $f$ is set to zero, and the domain $\Omega$ is defined as a circular region:
$\Omega=\{(x,y)\mid(x-0.5)^{2}+(y-0.5)^{2}\leq 0.5^{2}\}.$
The phantom coefficient $k$ within this domain is configured as shown in Figures 15(a) and 20(a). We conduct numerical experiments on two different phantom configurations of $k$, where a slight smoothing effect is applied using a Gaussian kernel. This reduces sharp transitions in the values and yields a more continuous distribution across the domain.
Since the inverse problem is ill-posed, we include a regularization term in the loss function to improve the stability and robustness of the reconstructed coefficient. Specifically, we extend the loss function (3) by introducing a total variation term [6], and use the following loss function:
$\mathfrak{L}(\tilde{k},\mathfrak{D},\tilde{\mathfrak{D}})=\sum_{i=1}^{N}\bigl{\|}h_{g_{i}}-\tilde{k}\,\partial_{\nu}\tilde{p}_{g_{i}}\bigr{\|}_{L^{2}(\partial\Omega)}^{2}+\lambda\bigl{\|}\nabla\tilde{k}\bigr{\|}_{L^{2}(\Omega)}^{2}.$ (6)
Here, $\mathfrak{D}=\{(g_{i},h_{g_{i}})_{i=1,\dots,N}\mid g_{i}\in H^{1/2}(\partial\Omega),h_{g_{i}}=k\partial_{\nu}p_{g_{i}}|_{\partial\Omega}\}$, $\tilde{\mathfrak{D}}=\{(g_{i},\tilde{k}\partial_{\nu}\tilde{p}_{g_{i}}|_{\partial\Omega})_{i=1,\dots,N}\mid g_{i}\in H^{1/2}(\partial\Omega)\}$. The positive parameter $\lambda$ acts as a regularization weight, controlling the contribution of the regularization term $\|\nabla\tilde{k}\|_{L^{2}(\Omega)}^{2}$.
The experimental setup for the phantom case is as follows. It is represented by a mesh with $N_{\text{phantom}}=51,433$ points. A Dirichlet boundary condition is imposed, given by $g=\cos(a\pi x)\cos(b\pi y)\big{|}_{\partial\Omega},$ where $a$ and $b$ are uniformly sampled from the interval $[0,2]$.
4.5.1 Example 3-1: Phantom 1
In this case, Phantom 1 is used as $k$ due to its relatively simple shape compared to the other phantom. We set the regularization weight in the loss function (6) as $\lambda=0.001/N_{\text{phantom}}$.
(a) exact $k$
(b) reconstructed $\tilde{k}$ using (3)
(c) absolute error $|k-\tilde{k}|$
(d) reconstructed $\tilde{k}$ using (6)
(e) absolute error $|k-\tilde{k}|$
Figure 12: Visualizations of the exact coefficient $k$, reconstructed coefficients $\tilde{k}$, and the corresponding absolute errors under two different loss functions. (c) shows the absolute error $|k-\tilde{k}|$ between (a) and (b), while (e) shows the absolute error between (a) and (d).
Figure 12 compares the results of the model trained with the regularization term included in the loss function (6) and the model trained without it using the loss function (3). The results show no significant difference between the two models, suggesting that the regularization term does not have a major impact due to the relative simplicity of the problem. Additionally, Figure 13 provides a three-dimensional illustration of the phantom configuration.
(a) exact $k$
(b) reconstructed $\tilde{k}$ using (3)
(c) reconstructed $\tilde{k}$ using (6)
Figure 13: Three-dimensional visualizations of the exact coefficient $k$ and the reconstructed coefficients $\tilde{k}$ obtained using the loss functions (3) and (6).
Figure 14: Relative error of $k$ with and without regularization term over epochs (log scale). Each point represents the averaged value over a 10 epoch interval.
Figure 14 plots the relative error of $k$ on a logarithmic scale. The results indicate that the model with the regularization term slightly outperforms the one without it, although the difference remains negligible. Both models exhibit a steady decrease in relative error and eventually converge to similar values. This suggests that the regularization term has little impact on the accuracy of the $k$ reconstruction in this case.
4.5.2 Example 3-1-2: Phantom 1 with Post-Processing
In this subsection, we extend the experiment conducted on Phantom 1 in Example 3-1 by incorporating a post-processing step. In this step, we apply the hyperbolic tangent function to constrain the reconstructed coefficient within the range $[0.2,1.0]$, assuming prior knowledge of this range. Apart from this post-processing step, the experimental setup remains the same as Example 3-1. We set the regularization weight in the loss function (6) as $\lambda=0.01/N_{\text{phantom}}$.
(a) exact $k$
(b) reconstructed $\tilde{k}$ using (3)
(c) absolute error $|k-\tilde{k}|$
(d) reconstructed $\tilde{k}$  (6)
(e) absolute error $|k-\tilde{k}|$
Figure 15: Visualizations of the exact coefficient $k$, reconstructed coefficients $\tilde{k}$, and the corresponding absolute errors under two different loss functions. (c) shows the absolute error $|k-\tilde{k}|$ between (a) and (b), while (e) shows the absolute error between (a) and (d).
Figure 15 compares the results of models trained using the loss function with the regularization term (6) and without it (3). The comparison indicates that the inclusion of the regularization term does not significantly affect the final reconstruction, as both models produce similar outcomes. However, a comparison between the results without post-processing (Figure 12) and with post-processing (Figure 15) shows that the region with noticeable error is significantly reduced in the post-processed case. This suggests that the post-processing step plays an important role in improving the spatial accuracy of the reconstructed coefficient.
Figure 16: Relative error of $k$ with and without regularization terms over epochs (log scale). Each point represents the averaged value over a 10 epoch interval.
Figure 16 presents the relative error in the coefficient $k$ for models trained with and without the regularization term, after applying the post-processing step. As in the case without post-processing, the model trained with the regularization term shows slightly better performance, but the difference remains negligible. This confirms that the regularization term has minimal impact on the accuracy of $k$ reconstruction, even after post-processing.
4.5.3 Example 3-2: Phantom 2
In this case, a more complex phantom, Phantom 2, is used to more clearly assess the effect of the regularization term. We set the regularization weight in the loss function (6) as $\lambda=0.1/N_{\text{phantom}}$.
(a) exact $k$
(b) reconstructed $\tilde{k}$ using (3)
(c) absolute error $|k-\tilde{k}|$
(d) reconstructed $\tilde{k}$ using (6)
(e) absolute error $|k-\tilde{k}|$
Figure 17: Visualizations of the exact coefficient $k$, reconstructed coefficients $\tilde{k}$, and the corresponding absolute errors under two different loss functions. (c) shows the absolute error $|k-\tilde{k}|$ between (a) and (b), while (e) shows the absolute error between (a) and (d).
Figure 17 presents visualizations of the exact coefficient $k$, the reconstructed coefficients $\tilde{k}$ obtained using the loss functions (3) and (6), and their corresponding absolute errors. This comparison illustrates the differences between the models trained with and without the regularization term.
(a) exact $k$
(b) reconstructed $\tilde{k}$ using (3)
(c) reconstructed $\tilde{k}$ using (6)
Figure 18: Three-dimensional visualizations of the exact coefficient $k$ and the reconstructed coefficients $\tilde{k}$ obtained using the loss functions (3) and (6).
To complement this, Figure 18 provides a three-dimensional visualizations of the same results. The overall structure of $k$ is well captured, but the reconstruction tends to overshoot at the peaks, producing sharper and taller spikes than the exact coefficient.
Figure 19: Relative error of $k$ with and without regularization terms over epochs (log scale). Each point represents the averaged value over a 10 epoch interval.
Figure 19 shows the relative error of $k$ over epochs on a logarithmic scale. While both models show similar convergence during the initial epochs of training, the model without regularization shows a slight increase in error in the later epochs. In contrast, the model with regularization maintains stable convergence throughout, suggesting that the regularization term improves training stability.
4.5.4 Example 3-2-2: Phantom 2 with Post-Processing
In this subsection, we extend the experiment conducted on Phantom 2 in Example 3-2 by incorporating a post-processing step. In this step, we apply the hyperbolic tangent function to constrain the reconstructed coefficient within the range $[1.0,5.0]$, assuming prior knowledge of this range. Apart from this post-processing step, the experimental setup remains the same as Example 3-2. We set the regularization weight in the loss function (6) as $\lambda=0.1/N_{\text{phantom}}$.
(a) exact $k$
(b) reconstructed $\tilde{k}$ using (3)
(c) absolute error $|k-\tilde{k}|$
(d) reconstructed $\tilde{k}$ using (6)
(e) absolute error $|k-\tilde{k}|$
Figure 20: Visualizations of the exact coefficient $k$, reconstructed coefficients $\tilde{k}$, and the corresponding absolute errors under two different loss functions. (c) shows the absolute error $|k-\tilde{k}|$ between (a) and (b), while (e) shows the absolute error between (a) and (d).
Figure 20 compares the reconstruction results of two models: one trained using the loss function (3), and the other using the loss function (6), which includes a regularization term. Figures 20(c) and 20(e) show the corresponding absolute error distributions. In Figure 20(c), large errors are observed around regions with sharp transitions in $k$, while Figure 20(e) shows smaller errors in these regions. This indicates that the model trained with the regularization term is better at capturing rapid changes and interfaces in the coefficient. These results suggest that the regularization term plays an important role in preserving the structural consistency of $k$, especially in complex phantom configurations.
(a) exact $k$
(b) reconstructed $\tilde{k}$ using the loss function (3) with $\omega=\partial\Omega$
(c) reconstructed $\tilde{k}$ using the loss function (6)
Figure 21: Illustrations of the exact coefficient $k$ and the reconstructed coefficient $\tilde{k}$ in three dimensions.
To complement this, Figure 21 provides a three-dimensional visualizations of the same results. The three-dimensional visualizations provide a clearer perspective on the shape and height of the reconstructed peaks. While both models capture the overall structure of $k$, the model with the regularization term—shown in Figure 21(c)—achieves more accurate peak heights and smoother transitions. In contrast, the model without the regularization term, shown in Figure 21(b), shows slight deformation around regions with sharp transitions in $k$.
Figure 22: Relative error of $k$ with and without regularization terms over epochs (log scale). Each point represents the averaged value over a 10 epoch interval.
Figure 22 shows the relative error of $k$ over epochs, plotted on a logarithmic scale. The model with regularization (red curve) shows a continuously decreasing error, especially after 30 epochs, and finally reaches a significantly lower final error than the model without regularization (blue curve). This behavior confirms that the regularization term not only improves accuracy but also stabilizes the training process, preventing overfitting and helping the model converge more reliably in complex scenarios.
Case Post-Processing MSE (without regularization) MSE (with regularization)
Phantom 1 No 0.00161 0.00155
Yes 0.00031 0.00029
Phantom 2 No 0.26601 0.21033
Yes 0.05494 0.02955
Table 2: MSE for Phantom 1 and Phantom 2 with and without regularization and post-processing.
Table 2 summarizes the MSE for Phantom 1 and Phantom 2 under different combinations of regularization and post-processing. For both phantoms, post-processing significantly improves the reconstruction accuracy, reducing the MSE by up to an order of magnitude. While the inclusion of the regularization term alone has only a marginal effect in Phantom 1, it is more effective in Phantom 2, especially when combined with post-processing. These results suggest that both regularization and post-processing contribute to improving reconstruction accuracy, with their combined effect being particularly important in more complex cases such as Phantom 2.
4.6 Example 4: Discontinuous Coefficient
In this example, we consider a more challenging and interesting case than the previous one, where the coefficient $k$ is discontinuous. The coefficient is defined as
$k(x,y)=\left\{\begin{array}[]{rl}0.9&\text{ if }(x-0.5)^{2}+(y-0.5)^{2}<(0.25)^{2}\\ 0.5&\text{ otherwise}\end{array}\right.$
as illustrated in Figure 23(a).
We conduct experiments for both the fully and partially observed settings. For the fully observed setting, the computational domain, number of mesh points, and Dirichlet boundary condition are the same as in Example 1. The source function is defined as $f(x,y)=k(x,y)(a^{2}+b^{2})\pi^{2}\cos(a\pi x)\cos(b\pi y)$. For the partially observed setting, both the Dirichlet boundary condition and the source function are set to zero on the region $\Gamma_{0}$, while all other setups remain the same as in the fully observed setting. In the post-processing step, we apply the hyperbolic tangent function to constrain the reconstructed coefficient within the range $[0.4,1.0]$.
In this example, we train the model using the loss function (3), considering both $\omega=\partial\Omega$ and $\omega=\Gamma$.
(a) exact $k$
(b) reconstructed $\tilde{k}$ using (3) with $\omega=\partial\Omega$
(c) absolute error $|k-\tilde{k}|$
(d) reconstructed $\tilde{k}$ using (3) with $\omega=\Gamma$
(e) absolute error $|k-\tilde{k}|$
Figure 23: Visualizations of the exact coefficient $k$, the reconstructed coefficients $\tilde{k}$ in three dimensions, and the corresponding absolute errors as a heatmap, for two different observation settings. (c) shows the absolute error $|k-\tilde{k}|$ between (a) and (b), while (e) shows the error between (a) and (d).
Figure 23 shows the reconstruction results and corresponding absolute errors for both the fully observed ($\omega=\partial\Omega$) and the partially observed ($\omega=\Gamma$) settings. The absolute error maps in Figures 23(c) and 23(e) show that the reconstruction errors are primarily concentrated around regions where $k$ is discontinuous. This suggests that the model struggles to capture sharp transitions, which is expected given the inherent difficulty of learning discontinuous coefficients.
(a) $y=x$
Figure 24: Slices of the exact and reconstructed coefficients along the diagonal line $y=x$.
Figure 24 shows a slice of the reconstructed coefficient $\tilde{k}$ along the diagonal line $y=x$. The blue dashed line represents the reconstruction from the fully observed setting ($\omega=\partial\Omega$), while the red dashed line corresponds to the reconstruction from the partially observed setting ($\omega=\Gamma$). The orange solid line indicates the exact coefficient $k$. Although small errors are observed near the discontinuities, both reconstructions closely approximate the true coefficient with minimal deviation.
Case Observed MSE MAE PSNR SSIM
Discontinuous Fully $2.28\times 10^{-4}$ $1.72\times 10^{-3}$ 35.51 0.9835
Partially $1.69\times 10^{-4}$ $1.11\times 10^{-3}$ 36.79 0.9896
Table 3: Comparison of MSE, MAE, PSNR, and SSIM for the discontinuous case under fully and partially observed settings.
Table 3 summarizes the MSE for the discontinuous case under both fully and partially observed settings. Similarly to the result for the spatially varying case, the MSE for the partially observed setting is slightly lower than that for the fully observed setting. This may be due to the fact that the reconstruction errors are mainly concentrated near the discontinuities of $k$, and both models achieve sufficiently accurate solutions in general. In such regimes, minor numerical differences during training may result in the partially observed model yielding a slightly better final reconstruction.
Acknowledgement
This work by D. Park and S. Moon are supported by the National Research Foundation of Korea, South Korea grant funded by the Korea government (NRF-2022R1C1C1003464 and RS-2023-00217116). The author S. Lee is supported by U.S. Department of Energy, Office of Science, Energy Earthshots Initiatives under Award DE-SC-0024703.

### v3

@@
 Inverse problems for Partial Differential Equations (PDEs) are crucial in numerous applications such as geophysics, biomedical imaging, and material science, where unknown physical properties must be inferred from indirect measurements.
-In this work, we address the inverse problem for elliptic PDEs by leveraging the Dirichlet-to-Neumann (DtN) map, which captures the relationship between boundary inputs and flux responses.
-Thus, this approach enables to solve the inverse problem that seeks the material properties inside the domain by utilizing the boundary data.
-Our framework employs an unsupervised machine learning algorithm that integrates a finite element method (FEM) in the inner loop for the forward problem, ensuring high accuracy.
-Moreover our approach is flexible to utilize partial observations of the boundary data, which is often the case in real-world scenarios.
-By incorporating carefully designed loss functions that accommodate discontinuities, the method refines coefficient reconstructions iteratively.
+In this work, we present a new approach to solving the inverse problem for elliptic PDEs, using only boundary data.
+Our method leverages the Dirichlet-to-Neumann (DtN) map, which captures the relationship between boundary inputs and flux responses.
+This enables the reconstruction of the unknown physical properties within the domain from boundary measurements alone.
+Our framework employs a self-supervised machine learning algorithm that integrates a Finite Element Method (FEM) in the inner loop for the forward problem, ensuring high accuracy.
+Moreover, our approach illustrates its effectiveness in challenging scenarios with only partial boundary observations, which is often the case in real-world scenarios.
+In addition, the proposed algorithm effectively handles discontinuities by incorporating carefully designed loss functions.
 This combined FEM and machine learning approach offers a robust, accurate solution strategy for a broad range of inverse problems, enabling improved estimation of critical parameters in applications from medical diagnostics to subsurface exploration.
 keywords:
-inverse problem, elliptic PDEs, EIT, variable coefficient, boundary condition mapping
+FEM, elliptic PDEs, inverse problem, Dirichlet-to-Neumann map, partial boundary observation
+journal: Journal of Computational and Applied Mathematics
 […]
@@
 1 Introduction
-Inverse problems for partial differential equations (PDEs) are a fundamental area of study in applied and computational mathematics, with wide-ranging applications in fields such as geophysics [1], biomedical imaging [2, 3], and material science.
-These problems center on deducing unknown coefficients—often representing physical properties such as conductivity, diffusivity, or elasticity—from observed data [4, 5].
-For instance, in medical imaging techniques such as Electrical Impedance Tomography (EIT), recovering these hidden material properties from indirect measurements enables visualization of conductivity distributions inside the human body, thereby enhancing diagnostic capabilities [6, 7, 8].
-Similarly, in geophysical contexts, inferring subsurface properties such as permeability from surface measurements is critical for resource exploration and environmental monitoring [9, 10].
-To address these challenges, we focus on the inverse problem for an elliptic equation with variable, potentially discontinuous coefficients.
-The main novelty of our approach is the utilization of the Dirichlet-to-Neumann (DtN) map—a mathematical construct that encapsulates the relationship between prescribed boundary inputs and their resulting flux responses [11].
+Inverse problems involve inferring unknown properties from indirect observations, in contrast to forward problems, where the governing equations and parameters are given and the solution is computed directly.
+Such problems often arise in science and engineering, particularly when reconstructing coefficients, sources, or initial conditions in governing equations.
+In particular, inverse problems for Partial Differential Equations (PDEs) are a fundamental area of study in applied and computational mathematics [Isakov06, Kirsch11], with wide-ranging applications in fields such as geophysics [Duchateau97], biomedical imaging [Kuchment13, Steinberg19], and material science [Janno11, Lee25].
+For instance, in medical imaging techniques such as Electrical Impedance Tomography (EIT), recovering these hidden material properties from indirect measurements enables visualization of conductivity distributions inside the human body, thereby enhancing diagnostic capabilities [Bar21, Cheney99, Patterson05].
+Similarly, in geophysical contexts, inferring subsurface properties such as permeability from surface measurements is critical for resource exploration and environmental monitoring [Ewing91, Carrera88].
+In recent years, machine learning techniques have been incorporated into the study of inverse problems, offering new perspectives beyond traditional approaches.
+Pioneering works have demonstrated that deep neural networks can solve PDE-related inverse problems directly, and improve reconstruction compared to classical methods [Jin17, Adler18].
+These developments have been systematically reviewed and categorized in comprehensive surveys on deep learning for inverse problems [Arridge19, Ongie20].
+Building on these advances, the neural operator framework was recently introduced to approximate solution operators between function spaces [Kovachki23], enabling the efficient resolution of parameterized PDE inverse problems.
+In parallel, the Finite Element Methods (FEM) has long played a central role in inverse problems, particularly in elasticity imaging.
+Early studies used the FEM to reconstruct material parameters from displacement data, and adjoint-based formulations later improved efficiency for biomedical applications [Maniatty89, Oberai03, Doyley12].
+FEM has also been applied to other inverse settings such as optical tomography [Arridge09], and recent work has extended its use to nonlinear constitutive models and hybrid approaches that integrate FEM with neural networks for efficient PDE inversion [Xu25, Meethal23].
+In this paper, we focus on the inverse problem for elliptic PDEs with variable, potentially discontinuous coefficients.
+The main novelty of our approach is the utilization of the Dirichlet-to-Neumann (DtN) map—a mathematical construct that encapsulates the relationship between prescribed boundary inputs and their resulting flux responses [Feldman19].
 This mapping framework is especially advantageous because it allows for the reconstruction of the interior characteristics of a domain using solely boundary measurements.
@@
 The presented numerical examples further illustrate the capabilities and robustness of the proposed algorithm, demonstrating its effectiveness across a range of challenging inverse problems.
-The proposed computational framework further distinguishes itself by integrating an unsupervised machine learning algorithm with finite element methods (FEM) used in the inner loop for solving the forward problem.
+The proposed computational framework further distinguishes itself by integrating a self-supervised machine learning algorithm [Subramanian23] with FEM used in the inner loop for solving the forward problem.
 This combination not only achieves high accuracy in modeling the physical processes but also enables iterative refinement of the coefficient reconstructions.
@@
 Ultimately, this framework offers a precise and reliable estimation of critical parameters, paving the way for significant advancements in applications ranging from non-invasive medical diagnostics to environmental and subsurface monitoring.
+In summary, our novelties include
+  1. 1.
+    Reconstruction of an unknown coefficient using only boundary measurements, without relying on interior data.
+  2. 2.
+    The framework demonstrates good performance even with partial boundary observations, reflecting realistic sensing scenarios.
+  3. 3.
+    It can reconstruct discontinuous coefficients by incorporating carefully designed loss functions.
+  4. 4.
+    A self-supervised framework is employed that learns from Cauchy data pairs, without requiring labeled interior coefficients.
+  5. 5.
+    The proposed method integrates the Finite Element Method (FEM) into the inner loop with adjoint-based gradients, utilizing FEM as a physics-grounded solver to ensure accuracy and consistency with the underlying PDE while maintaining computational efficiency in the inversion process.
 The rest of the paper is organized as follows.
 First, the governing system and the DtN mapping are introduced in Section 2.
-The overall computational framework including the loss functions is discussed in Section 3.
+The overall computational framework, including the loss functions, is discussed in Section 3.
 Finally, in Section 4, several numerical experiments to illustrate the capabilities and the performance of our algorithm are presented.
 2 Governing System
-In this paper, we consider the following equation
-$\left\{\begin{array}[]{rll}-\nabla\cdot(k({\mathbf{x}})\nabla p({\mathbf{x}}))&=f({\mathbf{x}})&{\mathbf{x}}\in\Omega,\\ p({\mathbf{x}})&=g({\mathbf{x}})&{\mathbf{x}}\in\partial\Omega,\end{array}\right.$ (1)
-where $\Omega$ is a bounded domain in $\mathbb{R}^{2}$, $p:\overline{\Omega}\rightarrow\mathbb{R}$ is the solution function, $f$ is the source function, $g\in H^{1/2}(\partial\Omega)$ is the given Dirichlet boundary condition, and $k$ is a strictly positive coefficient defined in $\Omega$.
+In this paper, we consider the following elliptic PDE
+$\left\{\begin{array}[]{rll}-\nabla\cdot(k({\mathbf{x}})\nabla p({\mathbf{x}}))&=f({\mathbf{x}})&{\mathbf{x}}\in\Omega,\\ p({\mathbf{x}})&=g({\mathbf{x}})&{\mathbf{x}}\in\partial\Omega,\end{array}\right$.
+(1)
+where $\Omega$ is a bounded domain in $\mathbb{R}^{2}$, $p:\overline{\Omega}\rightarrow\mathbb{R}$ is the solution function, $f$ is the source function, $g\in H^{1/2}(\partial\Omega)$ is the given Dirichlet boundary condition, and $k$ is a strictly positive coefficient defined in $\overline{\Omega}$.
+Here, ${\mathbf{x}}=(x,y)$ denotes the spatial coordinate in $\mathbb{R}^{2}$.
 We consider an inverse problem associated with (1), where the goal is to reconstruct the coefficient $k$ from the boundary measurements.
 To achieve this, we employ deep learning techniques.
-A straightforward approach would be to use supervised learning, which requires access to exact coefficient values of $k({\mathbf{x}})$ for all ${\mathbf{x}}\in\Omega$.
-However, in most cases, obtaining complete knowledge of $k$ in $\Omega$ is infeasible.
-Thus, we propose an unsupervised learning framework that does not rely on exact coefficient values of $k({\mathbf{x}})$ but instead leverages available data to infer the coefficient.
+A straightforward approach would be to use supervised learning, which requires access to exact coefficient values of $k({\mathbf{x}})$ for all ${\mathbf{x}}\in\overline{\Omega}$.
+However, in most cases, obtaining complete knowledge of $k$ in $\overline{\Omega}$ is infeasible.
+Thus, we propose a self-supervised learning framework that does not rely on exact coefficient values of $k$ but instead leverages available data such as Dirichlet and Neumann boundary conditions to infer the coefficient.
 The methodology and key concepts necessary for solving this problem are introduced in the following subsection.
 2.1 Dirichlet-to-Neumann Map
-In this subsection, we introduce the (Dirichlet-to-Neumann) DtN map, which is central to the reconstruction of $k$. This map is defined as follows:
+In this subsection, we introduce the DtN map, which is central to the reconstruction of $k$. This map is defined as follows:
 $\Lambda_{k}:H^{1/2}(\partial\Omega)\to H^{-1/2}(\partial\Omega),$
@@
 The space $H^{r}(\partial\Omega)$, $r\in\mathbb{R}$, denotes the Sobolev space.
-The reconstruction of the coefficient $k$ using the DtN map has been extensively studied in the context of Electrical Impedance Tomography (EIT) and the Calderón problem [11].
-In particular, the case when $f=0$ has been comprehensively surveyed in [12].
-Furthermore, additional studies have explored the uniqueness of $k$ under partially observed boundary data.
+The reconstruction of the coefficient $k$ using the DtN map has been extensively studied in the context of EIT and the Calderón problem [Feldman19].
+In particular, the case when $f=0$ has been comprehensively surveyed, including the uniqueness of $k$ under partially observed boundary data [Uhlmann14].
+Here, we briefly recapitulate the theorems that guarantee the existence and the uniqueness of $k$ with given conditions.
 Let $\Gamma\subset\partial\Omega$ be a non-empty open subset of the boundary.
-The following theorem proves the uniqueness of $k$.
-Theorem A (Corollary 1.19 in [12])
+Then the following theorem proves the uniqueness of $k$ in the case where $f=0$.
+Theorem A (Corollary 1.19 in [Uhlmann14])
 Let $\Omega\subset\mathbb{R}^{2}$ be a bounded domain with a smooth boundary, and let $k_{1},k_{2}\in C^{4+\alpha}(\overline{\Omega})$ be non-vanishing functions for some $\alpha>0$. Assume that
-$\Lambda_{k_{1}}(g)=\Lambda_{k_{2}}(g)\text{ on }\Gamma\text{ for all }g\in H^{1/2}(\partial\Omega),\text{ supp }g\subset\Gamma.$
+$\Lambda_{k_{1}}(g)=\Lambda_{k_{2}}(g)\text{ on }\Gamma\text{ for all }g\in H^{1/2}(\partial\Omega),\,\operatorname{supp}g\subset\Gamma.$
 Then, $k_{1}=k_{2}$ in $\Omega$.
 Furthermore, if $\Gamma=\partial\Omega$, the uniqueness of $k$ can be established using data from the entire boundary.
-Although the case $f\neq 0$ may seem easier because $f$ contains internal information, it remains an open problem.
-While the uniqueness of $k$ in the entire domain $\Omega$ is not guaranteed, its uniqueness on the boundary $\partial\Omega$ has been established [13].
-Theorem B (Theorem 5.1.1 in [13])
+Next, in the case where $f\neq 0$, we note that the uniqueness of $k$ in the entire domain $\Omega$ is not guaranteed.
+However, its uniqueness on the boundary $\partial\Omega$ has been established in [Isakov06] as the following theorem:
+Theorem B (Theorem 5.1.1 in [Isakov06])
 Assume that $k_{1},k_{2}\in C^{m}(V)$, where $V$ is a neighborhood of a boundary point of $\Omega$.
-[…]
-If $k_{1},k_{2}\in C(\Omega)$ are two coefficients that are analytic in $\Omega$ and $\Lambda_{k_{1}}=\Lambda_{k_{2}}$, then $k_{1}=k_{2}$ on $\partial\Omega$ and by Strong Maximum Principle [14, Chapter 2], we have $k_{1}=k_{2}$ in $\Omega$.
+ $\text{ If }\Lambda_{k_{1}}=\Lambda_{k_{2}},\text{ then }\partial^{\alpha}k_{1}=\partial^{\alpha}k_{2}\text{ on }\partial\Omega\cap V\text{ when }|\alpha|\leq m.$
+If $k_{1},k_{2}\in C(\Omega)$ are two coefficients that are analytic in $\Omega$ and $\Lambda_{k_{1}}=\Lambda_{k_{2}}$, then $k_{1}=k_{2}$ on $\partial\Omega$ and by Strong Maximum Principle [evans10, Chapter 2], we have $k_{1}=k_{2}$ in $\Omega$.
 Let $\omega$ be either $\Gamma$ or $\partial\Omega$.
-We refer to the case $\omega=\Gamma$ as the partially observed case, and $\omega=\partial\Omega$ as the fully observed setting.
+We refer to the case $\omega=\Gamma$ as the partially observed setting, and $\omega=\partial\Omega$ as the fully observed setting.
 Instead of directly using the DtN map, which is usually difficult to obtain in real-world scenarios, we utilize Cauchy data—pairs of Dirichlet and Neumann boundary measurements.
-$\left\{(g,h_{g}):g\in H^{1/2}(\partial\Omega),\text{ supp }g\subset\omega,h_{g}=k\partial_{\nu}p_{g}|_{\omega}\right\}.$
+$\left\{(g,h_{g}):g\in H^{1/2}(\partial\Omega),\,\operatorname{supp}g\subset\omega,h_{g}=k\partial_{\nu}p_{g}|_{\omega}\right\}.$
 These pairs serve as practical substitutes for the DtN map in our setup.
-Again, in practice, it is not feasible to train on all functions $g\in H^{1/2}(\partial\Omega)$ with supp $g\subset\omega$.
-We consider the following problem.
-Problem
-Given a sufficiently large finite collection of Cauchy data:
-$\left\{(g_{i},h_{g_{i}})_{i=1,\dots,N}:g_{i}\in H^{1/2}(\partial\Omega),\text{supp }g_{i}\subset\omega,h_{g_{i}}=k\partial_{\nu}p_{g_{i}}|_{\omega}\right\},$ (2)
+In practice, it is not feasible to train on all functions $g\in H^{1/2}(\partial\Omega)$ with supp $g\subset\omega$.
+We consider the following problem:
+Problem. Given a sufficiently large finite collection of Cauchy data:
+$\left\{(g_{i},h_{g_{i}})_{i=1,\dots,N}:g_{i}\in H^{1/2}(\partial\Omega),\,\operatorname{supp}g_{i}\subset\omega,h_{g_{i}}=k\partial_{\nu}p_{g_{i}}|_{\omega}\right\},$
 the coefficient $k$ can be numerically reconstructed in $\Omega$.
-Rather than directly using the values of $k$, we adopt an unsupervised learning framework where the model learns to reconstruct $k$ from Cauchy data data alone [15, 16].
-Using this approach, we reconstruct $k$ based only on Cauchy data, which is particularly useful in practical setups where $k$ is not directly measurable.
+Remark 1
+The above theorems establish the theoretical uniqueness of the coefficient $k$ under certain conditions on the DtN map.
+Based on these results, our work reconstructs $k$ from finitely many Cauchy data, as utilizing the full DtN map is not feasible in practice.
+Moreover, our framework handles both the fully observed setting ($\omega=\partial\Omega$) and the partially observed setting ($\omega=\Gamma$).
+Rather than directly using the values of $k$, we adopt a self-supervised learning framework where the model learns to reconstruct $k$ from Cauchy data alone [Hwang23, Hwang24].
+This approach is particularly useful in practical setups where $k$ is not directly measurable.
 3 Computational Algorithm
@@
 First, in Subsection 3.1, we provide an overview of the proposed framework.
-We introduce the model for reconstructing the coefficient $k$ in Subsection 3.2.
-Next, in Subsection 3.3, we describe how to solve the forward problem using the model described in 3.2.
+Then, we introduce the model for reconstructing the coefficient $k$ in Subsection 3.2.
+Next, in Subsection 3.3, we briefly describe how to solve the forward problem using the FEM.
 Finally, in Subsection 3.4, we define the loss function used to train the model and describe the method for calculating its gradient to update the model parameters.
@@
 In this subsection, we introduce the proposed computational framework for reconstructing the coefficient $k$.
-Input $\bullet$ Domain points
-$\overline{\Omega}$
-$\bullet$ Neumann data
-$h_{g}=k\partial_{\nu}p_{g}|_{\omega}$
-Neural Network $\mathcal{N}$
-$\bullet$ Multilayer Perceptron (MLP)
-: Computation results: Data flow Loss Function
-$\mathcal{L}\bigl{(}h_{g},\tilde{h}_{g}\bigr{)}=\bigl{\|}h_{g}(\cdot)-\tilde{h}_{g}(\cdot)\bigr{\|}^{2}_{L^{2}(\omega)}$
-Output $\bullet$ Reconstruct Coefficients
-$\{\tilde{k}({\mathbf{x}})\,|\,{\mathbf{x}}\!\!\in\!\!\overline{\Omega}\}$
-Predicted Neumann
-$\tilde{h}_{g}=\tilde{k}\,\partial_{\nu}\tilde{p}_{g}|_{\omega}$
-Solving for the PDE (FEniCS) Input $\bullet$ Source function $f$
-$\bullet$ Dirichlet condition $g$
-$\tilde{p}_{g}({\mathbf{x}})=\sum_{j=1}^{M}p_{j}\,\phi_{j}({\mathbf{x}})$ Parameter UpdatesDeep LearningFinite Element Method
+Input $\bullet$ Domain points  $\overline{\Omega}$ $\bullet$ Neumann data $h_{g}=k\partial_{\nu}p_{g}|_{\omega}$ Neural Network $\mathcal{N}$ $\bullet$ Multilayer Perceptron (MLP) : Computation results: Data flow Loss Function $\mathcal{L}\bigl(h_{g},\tilde{h}_{g}\bigr)=\bigl\|h_{g}(\cdot)-\tilde{h}_{g}(\cdot)\bigr\|^{2}_{L^{2}(\omega)}$ Output $\bullet$ Reconstruct Coefficient $\{\tilde{k}({\mathbf{x}})\,|\,{\mathbf{x}}\!\!\in\!\!\overline{\Omega}\}$ Predicted Neumann $\tilde{h}_{g}=\tilde{k}\,\partial_{\nu}\tilde{p}_{g}|_{\omega}$ Solving for the PDE (FEniCS) Input $\bullet$ Source function $f$ $\bullet$ Dirichlet condition $g$ $\tilde{p}_{g}({\mathbf{x}})=\sum_{j=1}^{M}p_{j}\,\phi_{j}({\mathbf{x}})$ Parameter Updates Deep Learning Finite Element Method
 Figure 2: Framework for reconstructing $k$ using a neural network and FEM.
 Figure 2 illustrates the process of reconstructing the coefficient $k$ using a neural network and FEM.
-The framework proceeds as follows: First, the domain points $\overline{\Omega}$ are input to the neural network $\mathcal{N}$, which outputs the reconstruction $\tilde{k}$ of $k$.
+The framework proceeds as follows: First, the domain points $\overline{\Omega}$ are the input to the neural network $\mathcal{N}$, which outputs the reconstruction $\tilde{k}$ of $k$.
 The neural network $\mathcal{N}$ is designed as a Multilayer Perceptron (MLP) that reconstructs $k$ from the input domain points.
-Next, the reconstructed coefficient $\tilde{k}$, along with the Dirichlet condition $g$ and source function $f$, is used to solve the (1) via FEM with the FEniCS package.
+Next, the reconstructed coefficient $\tilde{k}$, along with the Dirichlet condition $g$ and source function $f$, is used to solve (1).
 The resulting predicted solution is denoted as $\tilde{p}_{g}$, which is represented as a linear combination of basis functions $\{\phi_{i}\}_{i=1}^{M}$ in the finite-dimensional subspace $V_{h}$.
 These basis functions and the finite-dimensional subspace are further detailed in Subsection 3.3.
-Using $\tilde{k}$ and $\tilde{p}_{g}$, the predicted Neumann data $\tilde{k}\partial_{\nu}\tilde{p}_{g}|_{\omega}$ is computed.
-This predicted Neumann data is then compared with the given data $h_{g}$ to calculate the loss function.
+Using $\tilde{k}$ and $\tilde{p}_{g}$, the predicted Neumann data $\tilde{h}_{g}=\tilde{k}\partial_{\nu}\tilde{p}_{g}|_{\omega}$ is computed.
+This predicted Neumann data is then compared with the given Neumann data $h_{g}$ to calculate the loss function.
 Finally, the model parameters of the neural network $\mathcal{N}$ are updated to minimize the loss function.
@@
 In this subsection, we propose a model for reconstructing the coefficient $k$.
-The model is based on a MLP architecture, which leverages the Universal Approximation Theorem [17] as its theoretical foundation.
+The model is based on a MLP architecture, which leverages the universal approximation theorem [Cybenko89] as its theoretical foundation.
 This theorem ensures that neural networks, including MLPs, can approximate any continuous function to a desired level of accuracy.
@@
 Specifically, the reconstruction is expressed as $\tilde{k}({\mathbf{x}})=\mathcal{N}({\mathbf{x}};\boldsymbol{\theta})$, where $\boldsymbol{\theta}$ represents the learnable parameters of the model, such as weights and biases.
-The network consists of three hidden layers, each containing 50 hidden units.
-To enhance the ability of the model to capture periodic patterns and their derivatives, we employ the sine function as the activation function in the hidden layers.
-This choice enables the model to efficiently represent complex natural signals with high accuracy.
-3.3 Forward Problem
+The network consists of three hidden layers, each containing 50 hidden neurons.
+To enable the model to capture complex natural signals and their derivatives, we use the sine function as the activation function in the hidden layers.
+This periodic nonlinearity allows the model to represent fine details and high-frequency patterns more effectively than traditional activation functions [Sitzmann20].
+3.3 Finite Element Method for the Forward Problem
 In this subsection, we briefly describe the forward problem for the solution $p$ of (1) by utilizing FEM.
-Here, we employ the linear continuous Galerkin method, and it is implemented by FEniCS [18].
-The problem assumes a given Dirichlet condition $g$ and coefficient $k$, and the problem (1) is reformulated into a weak form by integration and then approximating the solution in a finite-dimensional space.
-In particular, let $V_{h}$ be a finite-dimensional subspace of $H^{1}(\Omega)$, and assume $p_{h},v_{h}\in V_{h}$, with $p_{h}$ representing the approximate solution and $v_{h}$ being the test function used in the weak formulation.
+Here, we employ the linear continuous Galerkin method, and it is implemented by FEniCS [Logg12].
+Given Dirichlet condition $g$ and coefficient $k$, (1) can be reformulated into a weak form by integration and then approximated in a finite-dimensional space.
+In particular, let $V_{h}$ be a finite-dimensional subspace of $H^{1}_{0}(\Omega)$ (to simplify the presentation, we assume $g=0$), and assume $p_{h},v_{h}\in V_{h}$, with $p_{h}$ representing the approximate solution and $v_{h}$ being the test function used in the weak formulation.
 Using the basis functions $\{\phi_{i}\}_{i=1}^{M}$ of $V_{h}$, the weak form can be approximated as:
-$\int_{\Omega}k\nabla p_{h}\cdot\nabla v_{h}\,{\rm d}{\mathbf{x}}=\int_{\Omega}fv_{h}\,{\rm d}{\mathbf{x}}.$
-Assuming $p_{h}=\sum_{i=1}^{M}P_{i}\phi_{i}$ and substituting $v_{h}=\phi_{j}$, the resulting system for each $j$-th term is:
-$\sum_{i=1}^{M}P_{i}\int_{\Omega}k\nabla\phi_{i}\cdot\nabla\phi_{j}\,{\rm d}{\mathbf{x}}=\int_{\Omega}f\phi_{j}\,{\rm d}{\mathbf{x}}.$
+$\int_{\Omega}k\nabla p_{h}\cdot\nabla v_{h}\,{\rm d}{\mathbf{x}}=\int_{\Omega}fv_{h}\,{\rm d}{\mathbf{x}},\ \forall v_{h}\in V_{h}.$
+Assuming $p_{h}=\sum_{i=1}^{M}P_{i}\phi_{i}$, where $P_{i}$ are constant coefficients, and substituting $v_{h}=\phi_{j}$, the resulting system becomes
+$\sum_{i=1}^{M}P_{i}\int_{\Omega}k\nabla\phi_{i}\cdot\nabla\phi_{j}\,{\rm d}{\mathbf{x}}=\int_{\Omega}f\phi_{j}\,{\rm d}{\mathbf{x}},\ \forall j=1,\cdots,M.$
 This can be written in matrix form as:
@@
 This subsection defines the loss function used to train the model introduced in Subsection 3.2, and describes how to compute its gradient with respect to the model parameters.
-The loss function based on the Problem is defined as
+The loss function based on Problem is defined as
 $\mathcal{L}(\mathcal{D},\tilde{\mathcal{D}})=\sum_{i=1}^{N}\bigl\|h_{g_{i}}-\tilde{k}\partial_{\nu}\tilde{p}_{g_{i}}\bigr\|_{L^{2}(\omega)}^{2},$ (2)
-where $\mathcal{D}=\{(g_{i},h_{g_{i}})_{i=1,\dots,N}\mid g_{i}\in H^{1/2}(\partial\Omega),\text{ supp }g_{i}\subset\omega,h_{g_{i}}=k\partial_{\nu}p_{g_{i}}|_{\omega}\}$, $\tilde{\mathcal{D}}=\{(g_{i},\tilde{k}\partial_{\nu}\tilde{p}_{g_{i}}|_{\omega})_{i=1,\dots,N}\mid g_{i}\in H^{1/2}(\partial\Omega),\text{ supp }g_{i}\subset\omega\}$, and $\|\cdot\|_{L^{2}(\omega)}$ denotes the $L^{2}$ norm on $\omega$.
+where $\mathcal{D}=\{(g_{i},h_{g_{i}})_{i=1,\dots,N}\mid g_{i}\in H^{1/2}(\partial\Omega),\,\operatorname{supp}g_{i}\subset\omega,h_{g_{i}}=k\partial_{\nu}p_{g_{i}}|_{\omega}\}$, $\tilde{\mathcal{D}}=\{(g_{i},\tilde{k}\partial_{\nu}\tilde{p}_{g_{i}}|_{\omega})_{i=1,\dots,N}\mid g_{i}\in H^{1/2}(\partial\Omega),\,\operatorname{supp}g_{i}\subset\omega\}$, and $\|\cdot\|_{L^{2}(\omega)}$ denotes the $L^{2}$ norm on $\omega$.
 The loss function $\mathcal{L}$ is minimized by adjusting the model parameters $\boldsymbol{\theta}$.
-When the loss function reaches its minimum, we have $\mathcal{D}=\tilde{\mathcal{D}}$.
+If the loss function reaches zero, then $\mathcal{D}=\tilde{\mathcal{D}}$.
 To optimize the model, we need the gradient of the loss function $\mathcal{L}$ with respect to the parameters $\boldsymbol{\theta}$.
 However, since this gradient cannot be computed directly, we apply the chain rule:
-$\frac{\partial\mathcal{L}}{\partial\boldsymbol{\theta}}=\frac{\partial\mathcal{L}}{\partial k_{\text{fem}}}\cdot\frac{\partial\tilde{k}}{\partial\boldsymbol{\theta}}.$ (4)
-[…]
+$\frac{\partial\mathcal{L}}{\partial\boldsymbol{\theta}}=\frac{\partial\mathcal{L}}{\partial\tilde{k}_{\text{fem}}}\cdot\frac{\partial\tilde{k}}{\partial\boldsymbol{\theta}}.$ (3)
+Here, $\tilde{k}$ represents the neural network output as a PyTorch tensor and $\tilde{k}_{\text{fem}}=\tilde{k}.\texttt{detach()}.\texttt{numpy()}$ is a coefficient that shares the same values as $\tilde{k}$ but is converted to a NumPy array for use in the forward PDE solve within FEniCS.
+The detach() operation disconnects $\tilde{k}$ from PyTorch’s autograd graph, meaning that any gradient information associated with it is lost beyond this point.
+As a result, the gradient $\partial\mathcal{L}/\partial\tilde{k}_{\text{fem}}$ is computed efficiently using the adjoint method in FEniCS, while $\partial\tilde{k}/\partial\boldsymbol{\theta}$ is obtained via PyTorch’s automatic differentiation [Paszke19].
+Multiplying these two terms yields the desired gradient $\partial\mathcal{L}/\partial\boldsymbol{\theta}$.
 4 Numerical Experiments
-In this section, we present numerical experiments to validate and demonstrate the effectiveness of the proposed framework introduced in Section 3.
+In this section, we present several numerical experiments to validate and demonstrate the effectiveness of the proposed framework introduced in Section 3.
 The following setup is common across all examples.
-The entire dataset consists of 4,096 samples: 2,048 for training, 1,024 for validation, and 1,024 for testing.
-The neural network is trained for 100 epochs with a batch size of 1.
-The learning rate is set to 0.001 for the first 50 epochs and reduced to 0.0001 for the remaining 50 epochs to facilitate stable convergence and prevent overshooting during the later stages of training.
-Training is performed using PyTorch [19] and the ADAM optimizer [20], a variant of stochastic gradient descent.
-All experiments are conducted on an NVIDIA GeForce RTX 3090 GPU (24GB VRAM), an Intel Core i9-12900K CPU, and 62GB RAM.
-The system runs with CUDA 11.8 and NVIDIA driver version 522.06.
-In the next subsections, we evaluate the performance of the proposed method in several cases:
-4.1 Example 1: Constant Coefficient
+The neural network is trained for 100 epochs with a batch size of 1, using PyTorch [Paszke19] and the ADAM optimizer [Kingma14], a variant of stochastic gradient descent.
+All experiments are performed on a system equipped with an NVIDIA GeForce RTX 3090 GPU (24GB VRAM), an Intel Core i9-12900K CPU, and 62GB of RAM.
+The software environment includes CUDA 11.8 and NVIDIA driver version 522.06.
+4.1 Example 1-1: Constant Coefficient
 In this example, we consider a constant coefficient $k$ to provide a baseline for evaluating model performance under simple conditions.
 Specifically, we set $k=1$.
-The computational domain is a unit square $\Omega=(0,1)\times(0,1)$ represented by a mesh with 66,049 points.
-The Dirichlet boundary condition is given by $g=\cos(a\pi x)\cos(b\pi y)|_{\partial\Omega}$, where, $a$ and $b$ are independently sampled from a uniform distribution over $[0,2]$.
-The source function is defined as $f(x,y)=(a^{2}+b^{2})\pi^{2}\cos(a\pi x)\cos(b\pi y)$.
-To evaluate how effectively our method reconstructs the coefficient $\tilde{k}$, we employ two measures: the Neumann data loss and the relative error of $k$.
+The computational domain is the unit square $\Omega=(0,1)\times(0,1)$, discretized using a triangular mesh with 4,225 degrees of freedom for the linear Lagrange finite element space.
+The Dirichlet boundary condition is defined as $g(x,y)=\cos(a\pi x)\cos(b\pi y)|_{\partial\Omega}$, where $a$ and $b$ are independently sampled from the uniform distribution over $[0,2]$.
+This experiment uses a Cauchy dataset containing 2,048 samples.
+The source function is given by $f(x,y)=-\nabla\cdot(k(x,y)\nabla(\cos(a\pi x)\cos(b\pi y)))$.
+The learning rate is initialized at 0.001 and reduced by a factor of 0.25 every 20 epochs to ensure stable convergence and prevent overshooting in the later stages of training.
+To evaluate the effectiveness of our method in reconstructing the coefficient $k$, we employ two measures: the Neumann data loss and the relative error of $\tilde{k}$.
 The Neumann data loss, defined in (2) with $\omega=\partial\Omega$ (fully observed setting), measures the difference between the predicted and observed Neumann data.
-Meanwhile, the relative error provides a normalized measure of the difference between $\tilde{k}$ and the exact coefficient $k$:
+Meanwhile, the relative error is defined as
 $\|\tilde{k}\|_{\text{rel}}=\frac{\|k-\tilde{k}\|_{L^{2}(\overline{\Omega})}}{\|k\|_{L^{2}(\overline{\Omega})}},$ (4)
-where $\|\cdot\|_{L^{2}(\overline{\Omega})}$ denotes the $L^{2}$ norm in $\overline{\Omega}$.
-Figure 3 presents the Neumann data loss (left) and the relative error of $k$ (right) side by side on a logarithmic scale, where each point represents the averaged value over a 10 epoch interval.
-This averaging smooths out short-term fluctuations, providing a clearer view of the overall training dynamics.
-Figure 3: (Left) Neumann training and validation loss (3) with $\omega=\partial\Omega$, and (Right) relative error of $k$ over epochs (log scale).
-Each point represents the averaged value over a 10 epoch interval.
-As shown in the left plot of Figure 3, the training loss (blue dashed line) decreases sharply at the beginning of the training process and continues to decrease, reaching around $10^{-4}$ during 20–50 epochs.
-After 50 epoch, when the learning rate is reduced, the training loss continues to decrease and finally stabilizes around $10^{-6}$.
-The validation loss (red line) shows larger oscillations due to the smaller validation set size and the fact that it is computed once per training epoch, which increases its sensitivity to parameter updates.
-Despite this variability, the validation loss also experiences a significant decrease after 50 epoch, reaching approximately $10^{-8}$.
-These observations suggest that the model makes effective use of the Neumann boundary data, with the learning rate schedule playing a crucial role in refining the parameter updates.
-On the right, the relative error of $k$ follows a similar trend as the Neumann training loss (blue dashed line on the left).
-After a rapid initial decrease, it remains around $10^{-3}$ during 20-50 epochs, then undergoes a sharp decrease and stabilizes approximately $10^{-4}$.
-This simultaneous decrease in both metrics indicates that reducing the Neumann data loss effectively guides $\tilde{k}$ to the exact coefficient $k$.
-These results support the claim in Problem that $k$ can be uniquely determined from Cauchy data.
-(a) exact $k$
-(b) reconstructed $\tilde{k}$
-(c) absolute error $|k-\tilde{k}|$
-Figure 4: Visualizations of the exact coefficient $k$, the reconstructed coefficient $\tilde{k}$, and the absolute error $|k-\tilde{k}|$.
-Figure 4 visualizes the exact coefficient $k$, the reconstructed coefficient $\tilde{k}$, and the absolute error $|k-\tilde{k}|$, providing a clear assessment of the reconstruction accuracy.
-The results show that the proposed method yields a highly accurate approximation of $k$ with only minor deviations.
-This confirms the effectiveness of the approach in reconstructing constant coefficients with minimal error.
-Up to this point, our analysis of the Neumann data loss and reconstruction error has demonstrated that the model can accurately reconstruct a constant coefficient.
-These numerical results confirm that training based on Neumann data plays a crucial role in optimizing the model.
+where $\|\cdot\|_{L^{2}(\overline{\Omega})}$ denotes the $L^{2}$ norm in $\overline{\Omega}$, and provides a normalized measure of the difference between $\tilde{k}$ and the exact coefficient $k$.
+(a) Neumann data loss and relative error of $\tilde{k}$
+(b) absolute error |kk~||k-$\tilde${k}|.
+Figure 3: Example 1-1.
+Training dynamics and reconstruction accuracy.
+(a) Neumann data loss (blue, left axis) and relative error of $\tilde{k}$ (red, right axis) over training epochs (log scale).
+Each point represents the averaged value over a 10-epoch interval.
+(b) Absolute error distribution between the exact coefficient $k$ and the reconstructed coefficient $\tilde{k}$.
+Figure 3 illustrates both the training dynamics and the reconstruction accuracy.
+Figure 3(a) shows the Neumann data loss (blue, left axis) and the relative error of $\tilde{k}$ (red, right axis), both plotted on a logarithmic scale.
+Each point corresponds to the average over a 10-epoch interval, which smooths out short-term fluctuations and highlights the overall decreasing trend.
+The simultaneous reduction of both metrics demonstrates that minimizing the Neumann data loss effectively drives $\tilde{k}$ toward the exact coefficient $k$, supporting the uniqueness claim in Problem.
+Figure 3(b) visualizes the absolute error |kk~||k-$\tilde${k}| between the exact coefficient $k$ and the reconstructed coefficient $\tilde{k}$.
+The results confirm that the proposed method provides a highly accurate approximation of $k$, with only minor deviations.
 To provide a more intuitive understanding of how the boundary conditions operate in practice, we now examine the operation of the DtN map as illustrated in Figure 4.
 […]
-Figure 5: Illustrations of the Dirichlet-to-Neumann (DtN) map $\Lambda$ for a 2D problem.
-The solution $p(x,y)$ is shown as a heat map.
-The Dirichlet boundary values $p(x,1)$ and $p(1,y)$ are mapped to their corresponding Neumann data via the DtN map.
+Figure 4: Example 1-1.
+Illustration of the Dirichlet-to-Neumann (DtN) map.
+The solution $p(x,y)$ is shown and the Dirichlet boundary values $p(x,1)$ and $p(1,y)$ mapped to their corresponding Neumann data via the DtN map are presented.
 The heat map on the left represents the solution $p(x,y)=\cos(0.3\pi x)\cos(0.6\pi y)$ over a rectangular domain.
@@
 4.2 Example 1-2: Constant Coefficient with Partially Observed Data
-In this example, we address the inverse problem of the partially observed data.
+In this example, we address the inverse problem under the partially observed data.
 In many real-world scenarios, measurements may be unavailable in certain regions due to sensor limitations or other constraints.
 We consider the case of a constant coefficient $k=1$.
-The computational domain defined as the unit square $\Omega=(0,1)\times(0,1)$, represented by a mesh with 66,049 points.
-To simulate partial observations, we assume that measurements are unavailable in certain parts of the boundary, covering approximately 20% of $\partial\Omega$.
-Specifically, we define the region $\Gamma_{0}$, where measurements are unavailable, as follows: measurements are unavailable for $y\in[\frac{3}{5},\frac{4}{5}]$ on the left boundary ($x=0$), for $y\in[\frac{1}{5},\frac{2}{5}]$ on the right boundary ($x=1$), for $x\in[\frac{2}{5},\frac{3}{5}]$ on the bottom boundary ($y=0$), and for $x\in[\frac{3}{5},\frac{4}{5}]$ on the top boundary ($y=1$).
+All other configurations, including the computational domain, degrees of freedom, the number of data samples, and the learning rate, are the same as in Example 1-1.
+To simulate partial observations, we assume that measurements are unavailable in certain parts of the boundary, covering approximately 40% of $\partial\Omega$.
+Specifically, we define the region $\Gamma_{0}$, where measurements are unavailable, as follows: measurements are unavailable for $y\in[\frac{2}{5},\frac{4}{5}]$ on the left boundary ($x=0$), for $y\in[\frac{1}{5},\frac{3}{5}]$ on the right boundary ($x=1$), for $x\in[\frac{1.5}{5},\frac{3.5}{5}]$ on the bottom boundary ($y=0$), and for $x\in[\frac{2.5}{5},\frac{4.5}{5}]$ on the top boundary ($y=1$).
 In these regions, both the Dirichlet boundary condition and the source function are set to zero; that is, $g=0$ and $f=0$ on $\Gamma_{0}$.
@@
 On $\Gamma$, the Dirichlet boundary condition is defined as $g(x,y)=\cos(a\pi x)\cos(b\pi y)$, where $a$ and $b$ are independently sampled from a uniform distribution over $[0,2]$.
-Similarly, the source function is defined on $\Gamma$ as $f(x,y)=(a^{2}+b^{2})\pi^{2}\cos(a\pi x)\cos(b\pi y)$.
+Similarly, the source function is defined on $\Gamma$ as $f(x,y)=-\nabla\cdot(k(x,y)\nabla(\cos(a\pi x)\cos(b\pi y)))$.
 The Neumann data is measured only on $\Gamma$, and the Neumann data loss function (2) is defined with $\omega=\Gamma$.
-Figure 6: (Left) Boundary regions and (Right) Dirichlet boundary condition $g$.
+Figure 5: Example 1-2. (Left) Boundary regions and (Right) Dirichlet boundary condition $g$. The values for which $g=0$ indicate that no data is available.
 Figure 5 illustrates the boundary region (left) and the corresponding Dirichlet boundary condition $g$ (right).
-Figure 7: Relative error of $k$ over epochs (log scale) for both fully and partially observed settings. Each point represents the average over a 10-epoch interval.
+(a) Relative error of $\tilde{k}$ over epochs (log scale) for fully and partially observed settings.
+(b) absolute error |kk~||k-$\tilde${k}|.
+Figure 6: Example 1-2.
+Training results and reconstruction accuracy for both fully and partially observed settings.
+(a) Relative error curves of $\tilde{k}$ under different observation settings.
+(b) Absolute error |kk~||k-$\tilde${k}| in the partially observed case.
+Figure 6 illustrates both the training results and the reconstruction accuracy.
+Figure 6(a) shows the relative error of $\tilde{k}$ on a logarithmic scale, comparing the fully observed ($\omega=\partial\Omega$) and partially observed ($\omega=\Gamma$) settings.
+Although the relative error in the partially observed setting is slightly higher than that in the fully observed case, it steadily decreases to a low value, indicating effective reconstruction.
+Figure 6(b) visualizes the absolute error |kk~||k-$\tilde${k}| between the exact coefficient $k$ and the reconstructed coefficient $\tilde{k}$ in the partially observed setting.
+These results indicate that the proposed method can accurately reconstruct the coefficient even with partial boundary measurements, demonstrating its robustness in handling incomplete data.
+4.3 Example 2-1: Spatially Varying Coefficient
+In this example, we test the algorithm with a spatially varying function $k$. Specifically, we set
+$k(x,y)=0.9\sin\left(\frac{\pi}{3}(x+y+0.1)\right).$
+All other configurations, including the computational domain, degrees of freedom, Dirichlet boundary condition, number of data samples, and learning rate, are the same as in Example 1-1.
+The source function is defined as $f(x,y)=-\nabla\cdot(k(x,y)\nabla(\cos(a\pi x)\cos(b\pi y)))$ by computing the divergence of the flux associated with the given function.
 (a) exact $k$
 (b) reconstructed $\tilde{k}$
-(c) absolute error $|k-\tilde{k}|$
-Figure 8: Visualizations of the exact coefficient $k$, the reconstructed coefficient $\tilde{k}$, and the absolute error $|k-\tilde{k}|$ in the partially observed setting.
-Figure 7 shows the relative error of $k$ on a logarithmic scale, comparing the cases of fully observed ($\omega=\partial\Omega$) and partially observed ($\omega=\Gamma$).
-Despite the limited observations in the partially observed setting, the relative error shows a convergence trend similar to that of the fully observed setting.
-This suggests that the proposed method remains robust even when only partial observations are available, achieving comparable accuracy.
-Figure 8 presents the exact coefficient $k$, the reconstructed coefficient $\tilde{k}$, and the absolute error $|k-\tilde{k}|$, providing a visual assessment of the reconstruction performance in the partially observed setting.
-The results indicate that the proposed method accurately reconstructs the coefficient even with partial boundary measurements, demonstrating its robustness in handling incomplete data.
-4.3 Example 2: Spatially Varying Coefficient
-In this example, we test the algorithm with a spatially varying function $k$. In this case, we set
-$k(x,y)=0.9\sin\left(\frac{\pi}{3}(x+y+0.1)\right).$
-All the other setups including the computational domain, number of mesh points, and Dirichlet boundary condition, are the same as the previous Example 1.
-The source function is defined as $f(x,y)=\nabla\cdot(k(x,y)\nabla(\cos(a\pi x)\cos(b\pi y)))$ by computing the divergence of the flux associated with the given function.
-(a) exact $k$
-(b) reconstructed $\tilde{k}$
-(c) absolute error $|k-\tilde{k}|$
-Figure 9: Illustrations of the exact coefficient $k$, the reconstructed coefficient $\tilde{k}$ in three dimensions, and the absolute error $|k-\tilde{k}|$ in a heat map.
-Figure 9 illustrates the exact coefficient $k$, the reconstructed coefficient $\tilde{k}$ in a three-dimensional space, and their absolute error in a heat map, providing a comparative view of their structures.
+(c) absolute error |kk~||k-$\tilde${k}|
+Figure 7: Example 2-1.
+Visualizations of the exact coefficient $k$ and the reconstructed coefficient $\tilde{k}$ using 3D surface plots, along with the absolute error |kk~||k-$\tilde${k}| shown as a 2D heat map.
+Figure 7 illustrates the exact coefficient $k$, the reconstructed coefficient $\tilde{k}$, and their absolute error, providing a comparative view of spatial structures.
 The reconstructed coefficient $\tilde{k}$ closely matches the exact $k$, effectively capturing its spatial variations, with the absolute error remaining at a low level.
@@
 (f) $y=0.5$
-Figure 10: Slices of the coefficient $\tilde{k}$ for the case of $k=0.9\sin\left(\dfrac{\pi}{3}(x+y+0.1)\right)$.
-Figure 10 shows the stability of the reconstruction process by comparing both the boundary and interior regions.
+Figure 8: Example 2-1. Slices of the exact coefficient $k$ and the reconstructed coefficient $\tilde{k}$ along selected lines in the domain.
+Figure 8 shows the reconstruction results by comparing both the boundary and interior regions.
 Over the entire domain, the reconstructed coefficient $\tilde{k}$ closely follows the behavior of $k$, showing strong agreement both at the boundaries and the interior.
@@
 4.4 Example 2-2: Spatially Varying Coefficient with Partially Observed Data
-In this example we consider the case of spatially varying coefficients on the partially observed data.
+In this example, we consider the case with partially observed data.
 The coefficient $k$ is defined as $k(x,y)=0.9\sin(\frac{\pi}{3}(x+y+0.1))$.
-All other setups including the computational domain, number of mesh point, Dirichlet condition, and loss function are the same as the previous Example 1-2.
-The source function is defined as $f(x,y)=\nabla\cdot(k(x,y)\nabla(\cos(a\pi x)\cos(b\pi y)))$ by computing the divergence of the flux associated with the given function.
+All other configurations, including the computational domain, degrees of freedom, Dirichlet boundary condition, number of data samples, and learning rate, are the same as in Example 1-2.
+The source function is defined as $f(x,y)=-\nabla\cdot(k(x,y)\nabla(\cos(a\pi x)\cos(b\pi y)))$ by computing the divergence of the flux associated with the given function.
 This function is imposed on $\Gamma$, while $f$ is set to zero on $\Gamma_{0}$.
+Figure 9: Example 2-2. Relative error of $\tilde{k}$ over epochs (log scale) for both fully and partially observed settings. Each point represents the average over a 10-epoch interval.
+Figure 9 shows the relative error of $\tilde{k}$ on a logarithmic scale, comparing the fully observed and partially observed settings.
+As in the constant coefficient case, the relative error under the partially observed setting is slightly higher, but it still decreases steadily, indicating good convergence behavior.
 (a) exact $k$
 (b) reconstructed $\tilde{k}$
-(c) absolute error $|k-\tilde{k}|$
-Figure 11: Visualizations of the exact coefficient $k$, the reconstructed coefficient $\tilde{k}$ in three dimensions, and the absolute error $|k-\tilde{k}|$ in a heat map in the partially observed setting.
-Figure 11 presents the exact coefficient $k$, the reconstructed coefficient $\tilde{k}$ in three-dimensional space, and the absolute error $|k-\tilde{k}|$ in a heat map, providing a visual comparison of the reconstruction performance under partially observed setting.
-The results confirm the robustness of the proposed method in accurately reconstructing the coefficient even with limited boundary measurements.
+(c) absolute error |kk~||k-$\tilde${k}|
+Figure 10: Example 2-2.
+Visualizations of the exact coefficient $k$ and the reconstructed coefficient $\tilde{k}$ using 3D surface plots, along with the absolute error |kk~||k-$\tilde${k}| shown as a 2D heat map in the partially observed setting.
+Figure 10 presents the exact coefficient $k$, the reconstructed coefficient $\tilde{k}$, and the absolute error |kk~||k-$\tilde${k}|, providing a visual comparison of the reconstruction performance under the partially observed setting.
+The results confirm the robustness of the proposed method in accurately reconstructing the spatially varying coefficient, even with limited boundary measurements.
 Case Observed MSE MAE PSNR SSIM
-Constant Fully $1.45\times 10^{-12}$ $1.05\times 10^{-6}$ 118.38 0.8346
-Partially $1.38\times 10^{-8}$ $1.00\times 10^{-3}$ 78.61 0.9438
-Spatially varying Fully $2.65\times 10^{-6}$ $1.50\times 10^{-3}$ 54.85 0.999996
-Partially $2.47\times 10^{-7}$ $3.80\times 10^{-4}$ 65.16 0.999997
-Table 1: Comparison of MSE, MAE, PSNR, and SSIM for the constant and spatially varying cases under fully and partially observed settings.
-Table 1 presents the mean squared errors (MSE) for the results of Examples 1, 1-2, 2, and 2-2, comparing the fully and partially observed settings for both constant and spatially varying cases.
-In the constant case, the model achieves extremely low reconstruction error, with an MSE of $1.45\times 10^{-12}$ in the fully observed setting.
-Even in the partially observed setting, the MSE remains low at $1.37\times 10^{-8}$, indicating strong reconstruction performance despite incomplete boundary data.
-In the spatially varying case, the reconstruction is more challenging due to the spatial variability of the coefficient $k$.
-Nonetheless, the model achieves a low MSE of $2.65\times 10^{-6}$ in the fully observed setting, demonstrating its capability to approximate the target function accurately.
-Interestingly, the MSE in the partially observed setting is lower than in the fully observed setting.
-This may be attributed to the fact that both models converge to highly accurate solutions with very small errors, where minor numerical variations can lead the partially observed setting to yield a slightly lower MSE.
+Constant Fully $3.10\times 10^{-11}$ $4.11\times 10^{-6}$ 105.09 0.999999999
+Partially $8.93\times 10^{-10}$ $2.11\times 10^{-5}$ 90.49 0.999999997
+Spatially varying Fully $95\times 10^{-8} 1.1$$11\times 10^{-4}$ 76.18 0.999999521
+Partially $4.10\times 10^{-8}$ $1.41\times 10^{-4}$ 72.95 0.999999575
+Table 1: Example 1-1 - Example 2-2. Comparison of MSE, MAE, PSNR, and SSIM for the constant and spatially varying cases under fully and partially observed settings.
+Table 1 presents the performance metrics—mean squared error (MSE), mean absolute error (MAE), peak signal-to-noise ratio (PSNR), and structural similarity index (SSIM)—for Examples 1-1, 1-2, 2-1, and 2-2.
+It compares the reconstruction quality under fully and partially observed settings for both the constant and spatially varying coefficient cases.
+PSNR measures the numerical similarity of the reconstructed values to the exact values [Huynh08], while SSIM evaluates the structural consistency between the reconstructed and exact values [Wang04].
+Higher values of PSNR and SSIM indicate better reconstruction performance, with SSIM values closer to 1 reflecting greater structural similarity.
+In contrast, lower values of MSE and MAE correspond to smaller reconstruction errors.
+In the constant case, the fully observed setting outperforms the partially observed setting in all metrics, including MSE, MAE, PSNR, and SSIM.
+Nevertheless, the partially observed setting achieves a very low MSE of $8.93\times 10^{-10}$ and a SSIM value close to 1, indicating that the overall reconstruction remains highly accurate.
+In the spatially varying case, all metrics are slightly lower than in the constant case, reflecting the increased complexity of the coefficient.
+The fully observed setting yields lower MSE and MAE values and achieves a higher PSNR, indicating a more accurate reconstruction.
+Although the SSIM value is slightly higher in the partially observed setting, the difference is negligible.
 4.5 Example 3: Phantom Coefficient with a Zero Source Function
-In this subsection, we consider the phantom coefficient, which is used in many medical imaging applications such as EIT [6, 7, 8] and PAT.
-In this case, we only consider the fully observed setting, where $\omega=\partial\Omega$. The source function $f$ is set to zero, and the domain $\Omega$ is defined as a circular region:
-$\Omega=\{(x,y)\mid(x-0.5)^{2}+(y-0.5)^{2}\leq 0.5^{2}\}.$
-The phantom coefficient $k$ within this domain is configured as shown in Figures 15(a) and 20(a).
-We conduct numerical experiments on two different phantom configurations of $k$, where a slight smoothing effect is applied using a Gaussian kernel.
+In this subsection, we consider the phantom coefficient, which is used in many medical imaging applications such as EIT [Bar21, Cheney99, Patterson05].
+In this case, we only consider the fully observed setting, where $\omega=\partial\Omega$.
+The source function $f$ is set to zero, and the domain $\Omega$ is defined as a circular region: $\Omega=\{(x,y)\mid(x-0.5)^{2}+(y-0.5)^{2}\leq 0.5^{2}\}.$ The phantom coefficient $k$ within this domain is configured as shown in Figures 11(a) and 16(a).
+We conduct numerical experiments on two different phantom coefficients $k$, where a slight smoothing effect is applied using a Gaussian kernel.
 This reduces sharp transitions in the values and yields a more continuous distribution across the domain.
-Since the inverse problem is ill-posed, we include a regularization term in the loss function to improve the stability and robustness of the reconstructed coefficient.
-Specifically, we extend the loss function (3) by introducing a total variation term [6], and use the following loss function:
+To enhance the accuracy in the case where the coefficient varies over the space, we include a regularization term in the loss function to improve the stability and robustness of the reconstructed coefficient [Bar21], and use the following loss function:
 $\mathfrak{L}(\tilde{k},\mathfrak{D},\tilde{\mathfrak{D}})=\sum_{i=1}^{N}\bigl\|h_{g_{i}}-\tilde{k}\,\partial_{\nu}\tilde{p}_{g_{i}}\bigr\|_{L^{2}(\partial\Omega)}^{2}+\lambda\bigl\|\nabla\tilde{k}\bigr\|_{L^{2}(\Omega)}^{2}.$ (5)
-Here, $\mathfrak{D}=\{(g_{i},h_{g_{i}})_{i=1,\dots,N}\mid g_{i}\in H^{1/2}(\partial\Omega),h_{g_{i}}=k\partial_{\nu}p_{g_{i}}|_{\partial\Omega}\}$, $\tilde{\mathfrak{D}}=\{(g_{i},\tilde{k}\partial_{\nu}\tilde{p}_{g_{i}}|_{\partial\Omega})_{i=1,\dots,N}\mid g_{i}\in H^{1/2}(\partial\Omega)\}$.
+Here, $\mathfrak{D}=\{(g_{i},h_{g_{i}})_{i=1,\dots,N}\mid g_{i}\in H^{1/2}(\partial\Omega),\,\operatorname{supp}g_{i}\subset\partial\Omega,h_{g_{i}}=k\partial_{\nu}p_{g_{i}}|_{\partial\Omega}\}$, $\tilde{\mathfrak{D}}=\{(g_{i},\tilde{k}\partial_{\nu}\tilde{p}_{g_{i}}|_{\partial\Omega})_{i=1,\dots,N}\mid g_{i}\in H^{1/2}(\partial\Omega),\,\operatorname{supp}g_{i}\subset\partial\Omega\}$.
 The positive parameter $\lambda$ acts as a regularization weight, controlling the contribution of the regularization term $\|\nabla\tilde{k}\|_{L^{2}(\Omega)}^{2}$.
 The experimental setup for the phantom case is as follows.
-It is represented by a mesh with $N_{\text{phantom}}=51,433$ points.
-A Dirichlet boundary condition is imposed, given by $g=\cos(a\pi x)\cos(b\pi y)\big{|}_{\partial\Omega},$ where $a$ and $b$ are uniformly sampled from the interval $[0,2]$.
-4.5.1 Example 3-1: Phantom 1
-In this case, Phantom 1 is used as $k$ due to its relatively simple shape compared to the other phantom.
-We set the regularization weight in the loss function (6) as $\lambda=0.001/N_{\text{phantom}}$.
+The Dirichlet boundary condition is imposed, given by $g(x,y)=\cos(a\pi x)\cos(b\pi y)\big|_{\partial\Omega},$ where $a$ and $b$ are uniformly sampled from the interval $[0,2]$.
+The circular domain is discretized using a triangular mesh with 51,433 degrees of freedom for the linear Lagrange finite element space.
+This experiment uses a Cauchy dataset containing 2,048 samples.
+The learning rate is set to 0.001 for the first 50 epochs and reduced to 0.0001 for the remaining 50 epochs.
+4.5.1 Example 3-1-1: Phantom 1
+In this case, Phantom 1 (Figure 11(a)) is used as the coefficient $k$ due to its relatively simple structure compared to the other phantom.
+The regularization weight in the loss function (5) is set to $\lambda=2\times 10^{-8}$.
+(a) Phantom 1: exact $k$
+(b) $\tilde{k}$ using (2)
+(c) absolute error |kk~||k-$\tilde${k}|
+(d) $\tilde{k}$ using (5)
+(e) absolute error |kk~||k-$\tilde${k}|
+Figure 11: Example 3-1-1.
+Phantom 1.
+Visualizations of the exact coefficient $k$, reconstructed coefficient $\tilde{k}$, and the corresponding absolute errors under two different loss functions.
+(c) shows the absolute error |kk~||k-$\tilde${k}| between (a) and (b), while (e) shows the absolute error between (a) and (d).
+Figure 11 compares the results of two models: one trained with the regularization term included in the loss function (5), and the other trained without it using the loss function (2).
+The results indicate that the inclusion of the regularization term does not lead to a significant difference in performance.
+Additionally, Figure 12 provides three-dimensional visualizations of the phantom coefficient.
 (a) exact $k$
-(b) reconstructed $\tilde{k}$ using (3)
-(c) absolute error $|k-\tilde{k}|$
-(d) reconstructed $\tilde{k}$ using (6)
-(e) absolute error $|k-\tilde{k}|$
-Figure 12: Visualizations of the exact coefficient $k$, reconstructed coefficients $\tilde{k}$, and the corresponding absolute errors under two different loss functions.
-(c) shows the absolute error $|k-\tilde{k}|$ between (a) and (b), while (e) shows the absolute error between (a) and (d).
-Figure 12 compares the results of the model trained with the regularization term included in the loss function (6) and the model trained without it using the loss function (3).
-The results show no significant difference between the two models, suggesting that the regularization term does not have a major impact due to the relative simplicity of the problem.
-Additionally, Figure 13 provides a three-dimensional illustration of the phantom configuration.
+(b) $\tilde{k}$ using (2)
+(c) $\tilde{k}$ using (5)
+Figure 12: Example 3-1-1. Three-dimensional visualizations of the exact coefficient $k$ and the reconstructed coefficient $\tilde{k}$ obtained using the loss functions (2) and (5).
+Figure 13: Example 3-1-1. Relative error of $\tilde{k}$ with and without regularization over epochs (log scale). Each point represents the averaged value over a 10-epoch interval.
+Figure 13 shows the relative error of $\tilde{k}$ on a logarithmic scale.
+The results indicate that the model trained with the regularization slightly outperforms the model trained without regularization, although the difference remains negligible.
+Both models exhibit a steady decrease in relative error and eventually converge to similar values.
+This suggests that, in this case, the regularization has little impact on the reconstruction accuracy of the coefficient $k$.
+4.5.2 Example 3-1-2: Phantom 1 with Post-Processing
+In this subsection, we extend the experiment conducted on Phantom 1 in Example 3-1-1 by incorporating a post-processing step.
+In this step, we apply the hyperbolic tangent function to constrain the reconstructed coefficient $\tilde{k}$ within the range $[0.2,1.0]$, based on the assumed prior knowledge of the range of $k$.
+The regularization weight in the loss function (5) is set to $\lambda=2\times 10^{-7}$.
+All other experimental configurations remain the same as in Example 3-1-1.
+(a) Phantom 1: exact $k$
+(b) $\tilde{k}$ using (2)
+(c) absolute error |kk~||k-$\tilde${k}|
+(d) $\tilde{k}$ using (5)
+(e) absolute error |kk~||k-$\tilde${k}|
+Figure 14: Example 3-1-2.
+Visualizations of the exact coefficient $k$, reconstructed coefficient $\tilde{k}$, and the corresponding absolute errors under two different loss functions with the proposed post-processing.
+(c) shows the absolute error |kk~||k-$\tilde${k}| between (a) and (b), while (e) shows the absolute error between (a) and (d).
+Figure 14 compares the results of models trained with and without the regularization term, using the loss functions (5) and (2), respectively.
+The comparison indicates that the inclusion of the regularization term does not significantly affect the final reconstruction, as both models produce similar outcomes.
+In contrast, a comparison between the results without post-processing (Figure 11) and with post-processing (Figure 14) shows a significant reduction in the region with noticeable error after post-processing.
+This suggests that the post-processing step plays an important role in improving the spatial accuracy of the reconstructed coefficient.
+Figure 15: Example 3-1-2.
+Relative error of $\tilde{k}$ with and without regularization terms, using the post-processing, over epochs (log scale).
+Each point represents the average over a 10-epoch interval.
+Figure 15 presents the relative error in the coefficient $k$ for models trained with and without the regularization term, after applying the post-processing step.
+As in the case without post-processing, the model trained with the regularization term shows slightly better performance, although the difference remains negligible.
+4.5.3 Example 3-2-1: Phantom 2
+In this case, a more complex phantom, Phantom 2 (Figure 16(a)), is used to more clearly assess the effect of the regularization term.
+The regularization weight in the loss function (5) is set to $\lambda=2\times 10^{-6}$.
+(a) Phantom 2: exact $k$
+(b) $\tilde{k}$ using (2)
+(c) absolute error |kk~||k-$\tilde${k}|
+(d) $\tilde{k}$ using (5)
+(e) absolute error |kk~||k-$\tilde${k}|
+Figure 16: Example 3-2-1.
+Phantom 2.
+Visualizations of the exact coefficient $k$, reconstructed coefficient $\tilde{k}$, and the corresponding absolute errors under two different loss functions.
+(c) shows the absolute error |kk~||k-$\tilde${k}| between (a) and (b), while (e) shows the absolute error between (a) and (d).
+Figure 16 presents visualizations of the exact coefficient $k$, the reconstructed coefficient $\tilde{k}$ obtained using the loss functions (2) and (5), and their corresponding absolute errors.
 (a) exact $k$
-(b) reconstructed $\tilde{k}$ using (3)
-(c) reconstructed $\tilde{k}$ using (6)
-Figure 13: Three-dimensional visualizations of the exact coefficient $k$ and the reconstructed coefficients $\tilde{k}$ obtained using the loss functions (3) and (6).
-Figure 14: Relative error of $k$ with and without regularization term over epochs (log scale). Each point represents the averaged value over a 10 epoch interval.
-Figure 14 plots the relative error of $k$ on a logarithmic scale.
-The results indicate that the model with the regularization term slightly outperforms the one without it, although the difference remains negligible.
-Both models exhibit a steady decrease in relative error and eventually converge to similar values.
-This suggests that the regularization term has little impact on the accuracy of the $k$ reconstruction in this case.
-4.5.2 Example 3-1-2: Phantom 1 with Post-Processing
-In this subsection, we extend the experiment conducted on Phantom 1 in Example 3-1 by incorporating a post-processing step.
-In this step, we apply the hyperbolic tangent function to constrain the reconstructed coefficient within the range $[0.2,1.0]$, assuming prior knowledge of this range.
-Apart from this post-processing step, the experimental setup remains the same as Example 3-1.
-We set the regularization weight in the loss function (6) as $\lambda=0.01/N_{\text{phantom}}$.
+(b) $\tilde{k}$ using (2)
+(c) $\tilde{k}$ using (5)
+Figure 17: Example 3-2-1. Three-dimensional visualizations of the exact coefficient $k$ and the reconstructed coefficient $\tilde{k}$ obtained using the loss functions (2) and (5).
+Furthermore, Figure 17 provides three-dimensional visualizations of the same results.
+The overall structure of $k$ is well captured, but the reconstruction tends to overshoot the peaks, producing sharper and taller spikes than the exact coefficient.
+Figure 18: Example 3-2-1. Relative error of $\tilde{k}$ with and without regularization terms over epochs (log scale). Each point represents the averaged value over a 10-epoch interval.
+Figure 18 shows the relative error of $\tilde{k}$ over epochs on a logarithmic scale.
+While both models show similar convergence in the initial training epochs, the model trained without regularization shows a slight increase in error in the later stages.
+In contrast, the model with regularization maintains stable convergence throughout, suggesting that the regularization term contributes to improved training stability.
+4.5.4 Example 3-2-2: Phantom 2 with Post-Processing
+In this subsection, we extend the experiment conducted on Phantom 2 in Example 3-2-1 by incorporating a post-processing step.
+In this step, we apply the hyperbolic tangent function to constrain the reconstructed coefficient $\tilde{k}$ within the range $[1.0,5.0]$, based on the assumed prior knowledge of the range of $k$.
+The regularization weight in the loss function (5) is set to $\lambda=2\times 10^{-6}$.
+All other experimental configurations remain the same as in Example 3-2-1.
+(a) Phantom 2: exact $k$
+(b) $\tilde{k}$ using (2)
+(c) absolute error |kk~||k-$\tilde${k}|
+(d) $\tilde{k}$ using (5)
+(e) absolute error |kk~||k-$\tilde${k}|
+Figure 19: Example 3-2-2.Visualizations of the exact coefficient $k$, reconstructed coefficient $\tilde{k}$, and the corresponding absolute errors under two different loss functions with post-processing.
+(c) shows the absolute error |kk~||k-$\tilde${k}| between (a) and (b), while (e) shows the absolute error between (a) and (d).
+Figure 19 compares the reconstruction results of two models: one trained using the loss function (2), and the other using the loss function (5), which includes a regularization term.
+Figures 19(c) and 19(e) show the corresponding absolute error distributions.
+In Figure 19(c), large errors are observed near regions with sharp transitions in $k$, while Figure 19(e) shows smaller errors in these regions.
+This indicates that the model trained with the regularization term is more effective at capturing rapid variations and interfaces in the coefficient.
+These results suggest that the regularization term plays a critical role in preserving the structural consistency of $k$, particularly in complex phantom coefficients.
 (a) exact $k$
-(b) reconstructed $\tilde{k}$ using (3)
-(c) absolute error $|k-\tilde{k}|$
-(d) reconstructed $\tilde{k}$  (6)
-(e) absolute error $|k-\tilde{k}|$
-Figure 15: Visualizations of the exact coefficient $k$, reconstructed coefficients $\tilde{k}$, and the corresponding absolute errors under two different loss functions.
-(c) shows the absolute error $|k-\tilde{k}|$ between (a) and (b), while (e) shows the absolute error between (a) and (d).
-Figure 15 compares the results of models trained using the loss function with the regularization term (6) and without it (3).
-The comparison indicates that the inclusion of the regularization term does not significantly affect the final reconstruction, as both models produce similar outcomes.
-However, a comparison between the results without post-processing (Figure 12) and with post-processing (Figure 15) shows that the region with noticeable error is significantly reduced in the post-processed case.
-This suggests that the post-processing step plays an important role in improving the spatial accuracy of the reconstructed coefficient.
-Figure 16: Relative error of $k$ with and without regularization terms over epochs (log scale). Each point represents the averaged value over a 10 epoch interval.
-Figure 16 presents the relative error in the coefficient $k$ for models trained with and without the regularization term, after applying the post-processing step.
-As in the case without post-processing, the model trained with the regularization term shows slightly better performance, but the difference remains negligible.
-This confirms that the regularization term has minimal impact on the accuracy of $k$ reconstruction, even after post-processing.
-4.5.3 Example 3-2: Phantom 2
-In this case, a more complex phantom, Phantom 2, is used to more clearly assess the effect of the regularization term.
-We set the regularization weight in the loss function (6) as $\lambda=0.1/N_{\text{phantom}}$.
-(a) exact $k$
-(b) reconstructed $\tilde{k}$ using (3)
-(c) absolute error $|k-\tilde{k}|$
-(d) reconstructed $\tilde{k}$ using (6)
-(e) absolute error $|k-\tilde{k}|$
-Figure 17: Visualizations of the exact coefficient $k$, reconstructed coefficients $\tilde{k}$, and the corresponding absolute errors under two different loss functions.
-(c) shows the absolute error $|k-\tilde{k}|$ between (a) and (b), while (e) shows the absolute error between (a) and (d).
-Figure 17 presents visualizations of the exact coefficient $k$, the reconstructed coefficients $\tilde{k}$ obtained using the loss functions (3) and (6), and their corresponding absolute errors.
-This comparison illustrates the differences between the models trained with and without the regularization term.
-(a) exact $k$
-(b) reconstructed $\tilde{k}$ using (3)
-(c) reconstructed $\tilde{k}$ using (6)
-Figure 18: Three-dimensional visualizations of the exact coefficient $k$ and the reconstructed coefficients $\tilde{k}$ obtained using the loss functions (3) and (6).
-To complement this, Figure 18 provides a three-dimensional visualizations of the same results.
-The overall structure of $k$ is well captured, but the reconstruction tends to overshoot at the peaks, producing sharper and taller spikes than the exact coefficient.
-Figure 19: Relative error of $k$ with and without regularization terms over epochs (log scale). Each point represents the averaged value over a 10 epoch interval.
-Figure 19 shows the relative error of $k$ over epochs on a logarithmic scale.
-While both models show similar convergence during the initial epochs of training, the model without regularization shows a slight increase in error in the later epochs.
-In contrast, the model with regularization maintains stable convergence throughout, suggesting that the regularization term improves training stability.
-4.5.4 Example 3-2-2: Phantom 2 with Post-Processing
-In this subsection, we extend the experiment conducted on Phantom 2 in Example 3-2 by incorporating a post-processing step.
-In this step, we apply the hyperbolic tangent function to constrain the reconstructed coefficient within the range $[1.0,5.0]$, assuming prior knowledge of this range.
-Apart from this post-processing step, the experimental setup remains the same as Example 3-2.
-We set the regularization weight in the loss function (6) as $\lambda=0.1/N_{\text{phantom}}$.
-(a) exact $k$
-(b) reconstructed $\tilde{k}$ using (3)
-(c) absolute error $|k-\tilde{k}|$
-(d) reconstructed $\tilde{k}$ using (6)
-(e) absolute error $|k-\tilde{k}|$
-Figure 20: Visualizations of the exact coefficient $k$, reconstructed coefficients $\tilde{k}$, and the corresponding absolute errors under two different loss functions.
-(c) shows the absolute error $|k-\tilde{k}|$ between (a) and (b), while (e) shows the absolute error between (a) and (d).
-Figure 20 compares the reconstruction results of two models: one trained using the loss function (3), and the other using the loss function (6), which includes a regularization term.
-Figures 20(c) and 20(e) show the corresponding absolute error distributions.
-In Figure 20(c), large errors are observed around regions with sharp transitions in $k$, while Figure 20(e) shows smaller errors in these regions.
-This indicates that the model trained with the regularization term is better at capturing rapid changes and interfaces in the coefficient.
-These results suggest that the regularization term plays an important role in preserving the structural consistency of $k$, especially in complex phantom configurations.
-(a) exact $k$
-(b) reconstructed $\tilde{k}$ using the loss function (3) with $\omega=\partial\Omega$
-(c) reconstructed $\tilde{k}$ using the loss function (6)
-Figure 21: Illustrations of the exact coefficient $k$ and the reconstructed coefficient $\tilde{k}$ in three dimensions.
-To complement this, Figure 21 provides a three-dimensional visualizations of the same results.
-The three-dimensional visualizations provide a clearer perspective on the shape and height of the reconstructed peaks.
+(b) $\tilde{k}$ using (2)
+(c) $\tilde{k}$ using (5)
+Figure 20: Example 3-2-2.
+Three-dimensional visualizations of the exact coefficient $k$ and the reconstructed coefficient $\tilde{k}$ obtained using the loss functions (2) and (5) with post-processing.
+To complement this, Figure 20 presents three-dimensional visualizations of the same results.
+These visualizations provide a clearer perspective on the shape and height of the reconstructed peaks.
 While both models capture the overall structure of $k$, the model with the regularization term—shown in Figure 20(c)—achieves more accurate peak heights and smoother transitions.
-In contrast, the model without the regularization term, shown in Figure 21(b), shows slight deformation around regions with sharp transitions in $k$.
-Figure 22: Relative error of $k$ with and without regularization terms over epochs (log scale). Each point represents the averaged value over a 10 epoch interval.
-Figure 22 shows the relative error of $k$ over epochs, plotted on a logarithmic scale.
-The model with regularization (red curve) shows a continuously decreasing error, especially after 30 epochs, and finally reaches a significantly lower final error than the model without regularization (blue curve).
-This behavior confirms that the regularization term not only improves accuracy but also stabilizes the training process, preventing overfitting and helping the model converge more reliably in complex scenarios.
-Case Post-Processing MSE (without regularization) MSE (with regularization)
-Phantom 1 No 0.00161 0.00155
-Yes 0.00031 0.00029
-Phantom 2 No 0.26601 0.21033
-Yes 0.05494 0.02955
-Table 2: MSE for Phantom 1 and Phantom 2 with and without regularization and post-processing.
-Table 2 summarizes the MSE for Phantom 1 and Phantom 2 under different combinations of regularization and post-processing.
-For both phantoms, post-processing significantly improves the reconstruction accuracy, reducing the MSE by up to an order of magnitude.
-While the inclusion of the regularization term alone has only a marginal effect in Phantom 1, it is more effective in Phantom 2, especially when combined with post-processing.
-These results suggest that both regularization and post-processing contribute to improving reconstruction accuracy, with their combined effect being particularly important in more complex cases such as Phantom 2.
+In contrast, the model without the regularization term, shown in Figure 20(b), shows slight deformation near regions with sharp transitions in $k$.
+Figure 21: Example 3-2-2.
+Relative error of $\tilde{k}$ with and without regularization terms, using the post-processing, over epochs (log scale).
+Each point represents the averaged value over a 10-epoch interval.
+Figure 21 shows the relative error of $\tilde{k}$ over epochs, plotted on a logarithmic scale.
+The model with regularization (red curve) shows a consistently decreasing error, especially after 30 epochs, and ultimately reaches a significantly lower final error than the model without regularization (blue curve).
+This behavior suggests that the regularization term not only improves reconstruction accuracy but also enhances training stability by promoting reliable convergence in complex scenarios.
+Case Post-Processing Regularization MSE MAE PSNR SSIM
+Phantom 1 No No 0.00162 0.02349 28.17 0.9575
+Yes 0.00155 0.02291 28.33 0.9583
+Yes No 0.00031 0.00721 35.03 0.9868
+Yes 0.00030 0.00708 35.30 0.9872
+Phantom 2 No No 0.26602 0.25938 23.66 0.9336
+Yes 0.21033 0.23995 24.20 0.9355
+Yes No 0.05494 0.11752 26.58 0.9488
+Yes 0.02955 0.09589 29.27 0.9653
+Table 2: Example 3.1-1 - 3.2-2. Comparison of metrics—MSE, MAE, PSNR, and SSIM—for Phantom 1 and Phantom 2 under different settings of post-processing and regularization.
+Table 2 summarizes the performance metrics—MSE, MAE, PSNR, and SSIM—for Phantom 1 and Phantom 2 under different combinations of post-processing and regularization.
+For both phantoms, post-processing leads to significant improvements in all metrics.
+These results suggest that while both regularization and post-processing improve reconstruction accuracy, their combined effect becomes particularly beneficial in more complex scenarios, such as Phantom 2.
 4.6 Example 4: Discontinuous Coefficient
-In this example, we consider a more challenging and interesting case than the previous one, where the coefficient $k$ is discontinuous. The coefficient is defined as
-$k(x,y)=\left\{\begin{array}[]{rl}0.9&\text{ if }(x-0.5)^{2}+(y-0.5)^{2}<(0.25)^{2}\\ 0.5&\text{ otherwise}\end{array}\right.$
+In this final example, we consider a more challenging and interesting case than the previous one, where the coefficient $k$ is discontinuous. The coefficient is defined as
+$k(x,y)=\left\{\begin{array}[]{rl}0.9&\text{ if }(x-0.5)^{2}+(y-0.5)^{2}<(0.25)^{2},\\ 0.5&\text{ otherwise},\end{array}\right$.
 as illustrated in Figure 22(a).
 We conduct experiments for both the fully and partially observed settings.
-For the fully observed setting, the computational domain, number of mesh points, and Dirichlet boundary condition are the same as in Example 1.
-The source function is defined as $f(x,y)=k(x,y)(a^{2}+b^{2})\pi^{2}\cos(a\pi x)\cos(b\pi y)$.
-For the partially observed setting, both the Dirichlet boundary condition and the source function are set to zero on the region $\Gamma_{0}$, while all other setups remain the same as in the fully observed setting.
-In the post-processing step, we apply the hyperbolic tangent function to constrain the reconstructed coefficient within the range $[0.4,1.0]$.
+All other configurations, including the computational domain, degrees of freedom, Dirichlet boundary condition, number of data samples, and learning rate, are the same as in Example 1-2.
+The source function is defined as $f(x,y)=-\nabla\cdot(k(x,y)\nabla(\cos(a\pi x)\cos(b\pi y)))$.
+For the partially observed setting, this source function is imposed on $\Gamma$, while $f$ is set to zero on $\Gamma_{0}$.
+We also include the post-processing step by applying the hyperbolic tangent function to constrain the reconstructed coefficient $\tilde{k}$ within the range $[0.4,1.0]$, based on the assumed prior knowledge of the range of $k$.
 In this example, we train the model using the loss function (2), considering both $\omega=\partial\Omega$ and $\omega=\Gamma$.
 (a) exact $k$
-(b) reconstructed $\tilde{k}$ using (3) with $\omega=\partial\Omega$
-(c) absolute error $|k-\tilde{k}|$
-(d) reconstructed $\tilde{k}$ using (3) with $\omega=\Gamma$
-(e) absolute error $|k-\tilde{k}|$
-Figure 23: Visualizations of the exact coefficient $k$, the reconstructed coefficients $\tilde{k}$ in three dimensions, and the corresponding absolute errors as a heatmap, for two different observation settings.
-(c) shows the absolute error $|k-\tilde{k}|$ between (a) and (b), while (e) shows the error between (a) and (d).
-Figure 23 shows the reconstruction results and corresponding absolute errors for both the fully observed ($\omega=\partial\Omega$) and the partially observed ($\omega=\Gamma$) settings.
-The absolute error maps in Figures 23(c) and 23(e) show that the reconstruction errors are primarily concentrated around regions where $k$ is discontinuous.
-This suggests that the model struggles to capture sharp transitions, which is expected given the inherent difficulty of learning discontinuous coefficients.
+(b) $\tilde{k}$ using (2) with $\omega=\partial\Omega$
+(c) absolute error |kk~||k-$\tilde${k}|
+(d) $\tilde{k}$ using (2) with $\omega=\Gamma$
+(e) absolute error |kk~||k-$\tilde${k}|
+Figure 22: Example 4.
+Visualizations of the exact coefficient $k$, the reconstructed coefficient $\tilde{k}$ in three dimensions, and the corresponding absolute errors as a heatmap, for two different observation settings.
+(c) shows the absolute error |kk~||k-$\tilde${k}| between (a) and (b), while (e) shows the error between (a) and (d).
+Figure 22 shows the reconstruction results and the corresponding absolute errors for both the fully observed and partially observed settings.
+The absolute error maps in Figures 22(c) and 22(e) show that the reconstruction errors are primarily concentrated near regions where $k$ is discontinuous.
+This indicates that the model has difficulty capturing sharp transitions, which is expected due to the inherent challenge of learning discontinuous coefficients.
 (a) $y=x$
-Figure 24: Slices of the exact and reconstructed coefficients along the diagonal line $y=x$.
-Figure 24 shows a slice of the reconstructed coefficient $\tilde{k}$ along the diagonal line $y=x$.
-The blue dashed line represents the reconstruction from the fully observed setting ($\omega=\partial\Omega$), while the red dashed line corresponds to the reconstruction from the partially observed setting ($\omega=\Gamma$).
-The orange solid line indicates the exact coefficient $k$.
-Although small errors are observed near the discontinuities, both reconstructions closely approximate the true coefficient with minimal deviation.
+Figure 23: Example 4. An illustration of the exact coefficient $k$ and the reconstructed coefficients $\tilde{k}$ under fully and partially observed settings, plotted along the diagonal line $y=x$.
+Figure 23 shows a slice of the exact coefficient $k$ and the reconstructed coefficients $\tilde{k}$ under the fully and partially observed settings along the diagonal line $y=x$.
+The solid orange line represents the exact coefficient, while the blue and red dashed lines correspond to the reconstructions from the fully and partially observed settings, respectively.
+Although slight differences appear near the discontinuities, both reconstructions closely follow the exact coefficient with minimal deviation.
 Case Observed MSE MAE PSNR SSIM
-Discontinuous Fully $2.28\times 10^{-4}$ $1.72\times 10^{-3}$ 35.51 0.9835
-Partially $1.69\times 10^{-4}$ $1.11\times 10^{-3}$ 36.79 0.9896
-Table 3: Comparison of MSE, MAE, PSNR, and SSIM for the discontinuous case under fully and partially observed settings.
-Table 3 summarizes the MSE for the discontinuous case under both fully and partially observed settings.
-Similarly to the result for the spatially varying case, the MSE for the partially observed setting is slightly lower than that for the fully observed setting.
-This may be due to the fact that the reconstruction errors are mainly concentrated near the discontinuities of $k$, and both models achieve sufficiently accurate solutions in general.
-In such regimes, minor numerical differences during training may result in the partially observed model yielding a slightly better final reconstruction.
+Discontinuous Fully $4.36\times 10^{-4}$ $2.81\times 10^{-3}$ 32.69 0.9733
+Partially $5.02\times 10^{-4}$ $3.30\times 10^{-3}$ 32.08 0.9676
+Table 3: Example 4. Comparison of MSE, MAE, PSNR, and SSIM for the discontinuous case under fully and partially observed settings.
+Table 3 presents the performance metrics—MSE, MAE, PSNR, and SSIM—for the discontinuous case under both the fully and partially observed settings.
+All metrics show slightly better performance in the fully observed setting, indicating a more accurate reconstruction overall.
+However, the partially observed setting also achieves highly competitive results.
+These findings suggest that the proposed method remains robust even with limited boundary data, particularly since reconstruction errors are primarily localized near discontinuities.
 Acknowledgement
-This work by D.
-Park and S.
-Moon are supported by the National Research Foundation of Korea, South Korea grant funded by the Korea government (NRF-2022R1C1C1003464 and RS-2023-00217116).
-The author S.
-Lee is supported by U.S.
+The work of D.
+Park was supported by the BK21 FOUR (Fostering Outstanding Universities for Research, No.
+202507810000), funded by the Ministry of Education (MOE, Korea) and the National Research Foundation of Korea (NRF-2022R1C1C1003464 and RS-2023-00217116).
+The work of S.
+Lee was supported by the U.S.
 Department of Energy, Office of Science, Energy Earthshots Initiatives under Award DE-SC-0024703.
+The work of S.
+Moon was supported by the National Research Foundation of Korea (NRF-2022R1C1C1003464 and RS-2023-00217116).